### v1

# A presentation of the pure cactus group of degree four

Abstract.
We give a simple presentation of the pure cactus group $PJ_{4}$ of degree four. This presentation is obtained by considering an action of $PJ_{4}$ on the hyperbolic plane and constructing a Dirichlet polygon for the action. As a corollary, we provide a direct alternative proof that $PJ_{4}$ is isomorphic to the fundamental group of the connected sum of five real projective planes.
Key words and phrases:
Cactus group, Cayley complex, Dirichlet polygon
2020 Mathematics Subject Classification:
20F65, 20F38, 05E18, 57M60
1. Introduction
As an analogue of the braid group, the cactus group $J_{n}$ was introduced in [9], motivated by the study of quantum groups. We will provide the purely algebraic definition used in this paper in the next section. It is worth noting that the group itself had been studied earlier in different contexts. See [4, 5] for examples.
The cactus group $J_{n}$ admits a natural surjection onto the symmetric group $S_{n}$ of degree $n$. The kernel of this surjection is called the pure cactus group $PJ_{n}$ of degree $n$. In [9, Theorem 9], Henriques and Kamnitzer showed that $PJ_{n}$ is isomorphic to the fundamental group of the Deligne–Mumford compactification $\overline{M_{0,n+1}}(\mathbb{R})$ of the moduli space of real genus-zero curves with $n+1$ marked points.
In the degree four case, it is known that $\overline{M_{0,5}}(\mathbb{R})$ is homeomorphic to the connected sum of five real projective planes (see [6, Example 2.5]). It follows that $PJ_{4}$ is isomorphic to the fundamental group of this surface, that is,
$\left\langle\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5}\mid\alpha_{1}^{2}\alpha_{2}^{2}\alpha_{3}^{2}\alpha_{4}^{2}\alpha_{5}^{2}=e\right\rangle.$
On the other hand, another explicit presentation of $PJ_{4}$ was obtained purely algebraically in [3, Appendix A], using the Reidemeister–Schreier method. This presentation is a simple one-relator presentation with five generators, but it is not related to the presentation described above. In fact, [3, Remark 5.6] states that it seems difficult to express the generators $\alpha_{k}$ explicitly in terms of the (standard) generators $s_{i,j}$ of the full cactus group $J_{4}$.
In this paper, we provide another presentation of the pure cactus group $PJ_{4}$ of degree four. This presentation is obtained geometrically by considering an action of $PJ_{4}$ on the hyperbolic plane and constructing a Dirichlet polygon for the action.
Theorem 1.1.
The pure cactus group $PJ_{4}$ admits the following presentation:
$\left\langle g_{1},\cdots,g_{10}\left|\begin{array}[]{l}g_{3}g_{6}^{-1}g_{7}g_{9}^{-1}g_{2}^{-1}=g_{3}g_{8}^{-1}g_{4}^{-1}=g_{5}g_{9}^{-1}g_{4}^{-1}\\ =g_{5}g_{1}g_{6}^{-1}=g_{8}g_{10}g_{7}^{-1}=g_{10}g_{1}^{-1}g_{2}=e\end{array}\right.\right\rangle$
This presentation is transformed to the next one:
$\left\langle g_{2},g_{4},g_{8},g_{9},g_{10}\mid g_{2}g_{9}g_{10}^{-1}g_{8}^{-1}g_{4}g_{9}g_{2}g_{10}g_{8}^{-1}g_{4}^{-1}=e\right\rangle$
We remark that our generators $g_{1},\dots,g_{10}$ are expressed as words in the standard generators of the cactus group $J_{4}$. In fact, the generators $g_{i}$ described above have minimal length with respect to the standard generating set $\{s_{i,j}\}$ of $J_{4}$; that is, their word lengths are either 4 or 5, and there are no elements of $PJ_{4}$ whose word length is 3 or less. See Section 4 for details.
It can be verified that the presentation given in Theorem 1.1 is equivalent to that of [3, Appendix A].
Corollary 1.2 ([3, Theorem 5.5]).
The pure cactus group $PJ_{4}$ of degree four admits the following presentation.
$\left\langle\alpha,\beta,\gamma,\delta,\epsilon\mid\alpha\gamma\epsilon\beta\epsilon\alpha^{-1}\delta^{-1}\beta\gamma\delta^{-1}=e\right\rangle$
As a corollary, we obtain a direct alternative proof of the fact that $PJ_{4}$ is isomorphic to the fundamental group of the connected sum of five real projective planes.
Corollary 1.3.
The pure cactus group $PJ_{4}$ of degree four admits the following presentation.
$\left\langle\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5}\mid\alpha_{1}^{2}\alpha_{2}^{2}\alpha_{3}^{2}\alpha_{4}^{2}\alpha_{5}^{2}=e\right\rangle.$
The key to the proof of the above theorem is an action of $PJ_{4}$ on the Cayley complex of the subgroup $J_{4}^{[2,3]}$ of $J_{4}$, which is isometric to the hyperbolic plane $\mathbb{H}^{2}$ up to scaling. We note that this action was originally introduced in [7].
Remark 1.
It is known that the moduli space $M_{0,n+1}(\mathbb{R})$ of real genus 0 curves over $\mathbb{R}$ with $n+1$ marked points can be identified with the configuration space of distinct $n+1$ points on the circle. See, for example, [4]. The idea of our proof in this paper comes from the cell complex structure of such configuration spaces studied in [1, 15, 10, 12]. Additionally, the relationship between the cactus group of degree 3 and such a configuration space is explored in [8].
2. Cactus group
In this section, we give definitions of the cactus group and the pure cactus group algebraically.
For a given integer $n\geq 2$, the cactus group of degree $n$, denoted by $J_{n}$, is defined as the group admitting a presentation with the generators $s_{p,q}$ for $1\leq p<q\leq n$ and the following relations.
  • $s_{p,q}^{2}=e$ for every $1\leq p<q\leq n$,
  • $s_{p,q}s_{m,r}=s_{m,r}s_{p,q}$ for all $1\leq p<q\leq n$ and $1\leq m<r\leq n$ satisfying $[p,q]\cap[m,r]=\emptyset$,
  • $s_{p,q}s_{m,r}=s_{p+q-r,p+q-m}s_{p,q}$ for all $1\leq p<q\leq n$ and $1\leq m<r\leq n$ satisfying $[m,r]\subset[p,q]$.
Here, $e$ denotes the trivial element of the group and $[p,q]$ denotes the set $\{p,p+1,\dots,q-1,q\}$ of integers for positive integers $p$, $q$ with $p<q$. Throughout the sequel, we denote the generators such as $s_{23}$ (without comma) in stead of $s_{2,3}$ for brevity.
For the degree four case, the cuctus group $J_{4}$ has the following presentation.
$\left\langle s_{12},s_{23},s_{34},s_{13},s_{24},s_{14}\left|\begin{array}[]{l}s_{12}^{2}=s_{23}^{2}=s_{34}^{2}=s_{13}^{2}=s_{24}^{2}=s_{14}^{2}=e,\\ s_{12}s_{34}=s_{34}s_{12},s_{12}s_{13}=s_{13}s_{23},s_{23}s_{24}=s_{24}s_{34},\\ s_{12}s_{14}=s_{14}s_{34},s_{23}s_{14}=s_{14}s_{23},s_{13}s_{14}=s_{14}s_{24}\end{array}\right.\right\rangle.$
Similar to the braid group, elements of the cactus group can be represented by diagrams of vertical strands on the plane. Some of such diagrams for $J_{4}$ are exhibited in Figure 1.
$s_{14}$$s_{13}$$s_{12}$$s_{23}$$s_{34}$
$=$$(s_{13})^{2}=e$$=$$s_{14}s_{12}=s_{34}s_{14}$
Figure 1. Diagrams for some elements of $J_{4}$
Due to this diagramatic expression, we see that the cactus group $J_{n}$ admits a natural projection $\pi:J_{n}\to S_{n}$ onto the symmetric group $S_{n}$ of degree $n$, and its kernel is called the pure cactus group of degree $n$, denoted by $PJ_{n}$. See [9, Subsection 3.1] or [7, Section 1] for more details.
We also introduce certain subgroup used in the sections later. For each integer $n\geq 2$ and subset $S\subset[2,n]$, let $J_{n}^{S}$ be the subgroup of $J_{n}$ generated by the elements $s_{p,q}$ for $1\leq p<q\leq n$ and $q-p+1\in S$ and defined by the following relations.
  • $s_{p,q}^{2}=e$ for every $1\leq p<q\leq n$ satisfying $q-p+1\in S$,
  • $s_{p,q}s_{m,r}=s_{m,r}s_{p,q}$ for every $1\leq p<q\leq n$ and $1\leq m<r\leq n$ satisfying $[p,q]\cap[m,r]=\emptyset$ and $q-p+1\in S$,
  • $s_{p,q}s_{m,r}=s_{p+q-r,p+q-m}s_{p,q}$ for every $1\leq p<q\leq n$ and $1\leq m<r\leq n$ satisfying $[m,r]\subset[p,q]$ and $q-p+1\in S$.
See [7, Section 5] for more details.
In later sections, we mainly consider the group $J_{4}^{[2,3]}$, which we denote by $J_{4}^{\prime}$ in the rest of paper. It is a subgroup of $J_{4}$ generated by $s_{12},s_{23},s_{34},s_{13},s_{24}$ (except for $s_{14}$), and it has the following presentation.
$\left\langle s_{12},s_{23},s_{34},s_{13},s_{24}\left|\begin{array}[]{l}s_{12}^{2}=s_{23}^{2}=s_{34}^{2}=s_{13}^{2}=s_{24}^{2}=e,\\ s_{12}s_{34}=s_{34}s_{12},s_{12}s_{13}=s_{13}s_{23},s_{23}s_{24}=s_{24}s_{34}\end{array}\right.\right\rangle.$
We denote by $\mathcal{C}$ the Cayley complex of the group $J_{4}^{\prime}$ with respect to the presentation above.
3. Preliminary
In this section, we prepare some properties of group action for the proof of the Theorem 1.1.
The following two propositions are given in [14]. See [14] for definitions of the terms used.
Proposition 3.1 ([14, Theorem 6.6.13]).
Let $D(a)$ be the Dirichlet domain with center $a$ for a discontinuous group $G$ of isometries of a metric space $X$ such that
  1. (1)
    $X$ is geodesically connected,
  2. (2)
    $X$ is geodesically complete,
  3. (3)
    $X$ is finitely compact.
Then $D(a)$ is a locally finite fundamental domain for $G$.
Proposition 3.2 ([14, Exercise 6.6, 2.]).
Let G be a group of isometries of a connected metric space $X$ with locally finite fundamental domain $R$. Then $G$ is generated by $\{g\in G\mid\overline{R}\cap g\overline{R}\neq\emptyset\}$.
Next we recall the following well-known theorem originated in [13]. Our terminologies follows from [11].
Proposition 3.3 (Poincaré’s polygon theorem, c.f. [11, p223]).
Let $D$ be a Poincaré polygon. Let $G$ be the group generated by the identifying generators. Then $G$ is discontinuous, $D$ is a fundamental polygon for $G$, and the cycle relations form the complete set of relations for $G$.
In the following, we give explanations for the terms used in the theorem above.
Let $D$ be a polygon on the hyperbolic plane $\mathbb{H}^{2}$. An identification on $D$ is a map which assigns, to each side $s$, a side $s^{\prime}$ and an isometry $A(s,s^{\prime})$ of $\mathbb{H}^{2}$ so that
  1. (i)
    $A(s,s^{\prime})$ maps $s$ onto $s^{\prime}$,
  2. (ii)
    $(s^{\prime})^{\prime}=s$ and $A(s^{\prime},s)=(A(s,s^{\prime}))^{-1}$,
  3. (iii)
    if $s=s^{\prime}$, then $A(s,s^{\prime})$ is the identity on $s$, and
  4. (iv)
    for each side $s$, there exists a neighborhood $V$ of $s$ so that $A(s,s^{\prime})(V\cap D)\cap D=\emptyset$.
We call the isometries $A(s,s^{\prime})$ which identify the sides of the polygon $D$ identifying generators, or generators.
Let $D$ be a polygon with identification. The identified polygon $D^{*}$ is the quotient obtained by the surjection $p:\overline{D}\longrightarrow D^{*}$ such that $p(x)=p(x^{\prime})$ if there exists a generator $A(s,s^{\prime})$ with $A(s,s^{\prime})(x)=x^{\prime}$. For two elements $x$, $x^{\prime}$ in $D^{*}$, we set a metric on $D^{*}$ as
$d_{D^{*}}^{*}(x,x^{\prime}):=\inf\sum_{i=e}^{n}d_{\mathbb{H}^{2}}(z_{i},z_{i}^{\prime})$
where the infimum is taken over all $n$ and over all $2n$-tuples of points of $\overline{D}$ such that $p(z_{1})=x$, $p(z_{i}^{\prime})=p(z_{i+1})$, and $p(z_{n}^{\prime})=x^{\prime}$.
Let $D$ be a polygon with identification. The polygon $D$ is complete if
  1. (v)
    for each $x\in D^{*}$, $p^{-1}(x)$ is a finite set,
  2. (vi)
    $D^{*}$ is complete in the metric $d^{*}_{\mathbb{H}^{2}}$.
Let $D$ be a complete polygon with identification. Let $z_{1}$ be some vertex of $D$. There are precisely two sides of $D$ which meet at $z_{1}$; we choose one of these and call it $s_{1}$. There is then a corresponding side $s_{1}^{\prime}$ and a generator $A_{1}=A(s_{1},s_{1}^{\prime})$. Set $z_{2}=A_{1}(z_{1})$ and observe that there is a unique other side $s_{2}$ which has $z_{2}$ as an end point. There is a corresponding side $s_{2}^{\prime}$ and a generator $A_{2}=A(s_{2},s_{2}^{\prime})$. Set $z_{3}=A_{2}(z_{2})$ and observe that there is a unique other side $s_{3}$ which has $z_{3}$ as an end point. In this manner, we get a sequence $\{z_{i}\}$ of vertices, and a sequence $\{A_{i}\}$ of generators. Then there exists periodic sequence of generators $\{A_{1},\cdots A_{n}\}$ such that $A_{n}(\cdots A_{2}(A_{1}(z_{1})))=z_{1}$ which is called a cycle of generators. For a cycle of generators $\{A_{1},\dots,A_{n}\}$, there exists a sequence of vertices $\{z_{1},\dots,z_{n}\}$ in $D$ such that $z_{2}=A_{1}(z_{1})$, …, $z_{n}=A_{n-1}(z_{n-1})$, and $A_{n}(z_{n})=z_{1}$. The sequence of the vertices is called a cycle of vertices. Then, the polygon $D$ is said to satisfy the cycle condition if for each cycle of vertices $\{z_{1},\dots,z_{n}\}$, there exists an integer $\nu$ so that
  1. (vii)
    $\nu\sum_{i=e}^{m}\alpha(z_{i})=2\pi$,
where $\alpha(z_{i})$ is the angle around $x_{i}$ measured from inside $D$.
Let $D$ be a polygon with identifications. The polygon $D$ is a Poincaré polygon if $D$ is complete and satisfies the cycle condition.
For a cycle of generators $\{A_{1},\dots,A_{n}\}$, by Condition (vii) , the $\nu$ times of the composition $A_{n}\circ\cdots\circ A_{1}$ is the identity on $D$, i.e., we get a relation $(A_{n}\circ\cdots\circ A_{1})^{\nu}=e$ for that cycle. We call such relations the cycle relation.
4. Proof of Theorem 1.1
Before proving Theorem 1.1, we prepare the following key proposition.
Proposition 4.1.
The Caylay complex $\mathcal{C}$ of $J_{4}^{\prime}$ is isometric to the hyperbolic plane $\mathbb{H}^{2}$ up to scaling.
Proof.
As shown in [7, Proposition 2.6], every element of the cactus group has a unique diagram of the normal form. This enable us to enumerate all the elements of $J_{4}^{\prime}$ with bounded word length, equivalently, to determine the local structure, vertices and edges, in $\mathcal{C}$ inside the bounded ball centered at $e$. Note that the vertices of $\mathcal{C}$ are identified with elements in $J_{4}^{\prime}$.
The vertices in $\mathcal{C}$ whose word length is $1$ are just generators;
$\begin{array}[]{ccccc}s_{12}&s_{23}&s_{34}&s_{13}&s_{24}.\end{array}$
The following are the vertices whose word length with respect to the generators above is $2$.
$\begin{array}[]{ccccc}s_{12}s_{23}&s_{13}s_{23}&s_{13}s_{24}&s_{13}s_{34}&s_{13}s_{12}\\ s_{23}s_{12}&s_{23}s_{34}&s_{24}s_{34}&s_{24}s_{12}&s_{24}s_{13}\\ s_{24}s_{23}&s_{34}s_{23}&s_{34}s_{13}&s_{34}s_{12}&s_{12}s_{24}\end{array}$
As shown Figure 2, There are $25$ edges ($1$-cells) connecting the above vertices (00-cells) in $\mathcal{C}$, and the $2$-cells (quadrangles) containing $e$ on $\mathcal{C}$ are
$\begin{array}[]{l}\langle e,s_{12},s_{12}s_{13},s_{13}\rangle,\langle e,s_{13},s_{13}s_{12},s_{23}\rangle,\langle e,s_{23},s_{23}s_{24},s_{24}\rangle,\\ \langle e,s_{24},s_{24}s_{23},s_{34}\rangle,\langle e,s_{34},s_{34}s_{12},s_{12}\rangle.\end{array}$
[…]
Figure 2. Neighborhood of $e$ in the Cayley complex $\mathcal{C}$
Therefore, $\mathcal{C}$ is a cell complex such that there are five quadrilateral $2$-cells surrounding each vertices.
We consider the map $f$ corresponding each closed $2$-cell on $\mathcal{C}$ to the square (regular quadrilateral) in $\mathbb{H}^{2}$ whose internal angles are equal to $\frac{2\pi}{5}$. The image of the map $f$ gives a uniform tiling (regular tessellation) of $\mathbb{H}^{2}$ of type $\{4,5\}$. See Figure 3.
Figure 3. The $\{4,5\}$-tesselation of the hyperbolic plane. ([2, Figure 7])
We see that this map is an isometry up to scaling $d(f(x),f(y))=Rd(x,y)$ with $R=\cosh^{-1}(\cot^{2}{\frac{\pi}{5}})$ by elementary calculation based on the hyperbolic trigonometry. ∎
In other words, if we set a metric on $\mathcal{C}$ where the length of each edge is the constant $R$, then $\mathcal{C}$ is truly isometric to $\mathbb{H}^{2}$. Therefore, we can apply results for the hyperbolic plane to the Cayley complex $\mathcal{C}$.
Now we consider an action of $PJ_{4}$ on $\mathcal{C}$. The following proposition was essentially obtained in [7].
Proposition 4.2 ([7, Proof of Corollary 7.3]).
The map $\Gamma$ induced from $\Gamma_{0}$ defined by the following implies an isometric action of $PJ_{4}$ on $\mathcal{C}$. Note that $\left(\mathcal{C}\right)^{(0)}$ is identified with $J_{4}^{\prime}$.
$\Gamma_{0}:PJ_{4}\times\left(\mathcal{C}\right)^{(0)}$ $\longrightarrow\left(\mathcal{C}\right)^{(0)}$
$(g,h)$$\longmapsto\begin{cases}gh&gh\in J_{4}^{\prime}\\ ghs_{14}&gh\notin J_{4}^{\prime}\end{cases}$
Moreover, the action $\Gamma$ on $\mathcal{C}$ is free and cocompact.
In the following, for simplicity, we identify $PJ_{4}$ and its image under the action $\Gamma$, and write $g\cdot x$ in stead of $\Gamma(g,x)$ for $g\in PJ_{4}$ and $x\in\mathcal{C}$. In this setting, $PJ_{4}$ is regarded as a subgroup of $\mathrm{Isom}\;\mathcal{C}$.
Now, let $a_{1},\cdots,a_{20}$ be the elements of $J_{4}^{\prime}$ defined as follows.
$\begin{array}[]{lll}a_{1}:=s_{13}s_{24}s_{12}s_{34}&a_{2}:=s_{13}s_{24}s_{13}s_{24}&a_{3}:=s_{13}s_{34}s_{23}s_{12}\\ a_{4}:=s_{13}s_{34}s_{13}s_{23}&a_{5}:=s_{23}s_{12}s_{23}s_{13}&a_{6}:=s_{23}s_{12}s_{24}s_{12}\\ a_{7}:=s_{23}s_{34}s_{13}s_{34}&a_{8}:=s_{24}s_{34}s_{23}s_{34}&a_{9}:=s_{24}s_{12}s_{24}s_{23}\\ a_{10}:=s_{24}s_{12}s_{23}s_{34}&a_{11}:=s_{24}s_{13}s_{24}s_{13}&a_{12}:=s_{24}s_{23}s_{13}s_{34}\\ a_{13}:=s_{34}s_{23}s_{34}s_{24}&a_{14}:=s_{34}s_{23}s_{12}s_{24}&a_{15}:=s_{34}s_{13}s_{34}s_{23}\\ a_{16}:=s_{34}s_{13}s_{23}s_{24}&a_{17}:=s_{34}s_{12}s_{24}s_{13}&a_{18}:=s_{12}s_{24}s_{12}s_{23}\\ a_{19}:=s_{12}s_{23}s_{34}s_{13}&a_{20}:=s_{12}s_{23}s_{12}s_{13}&\end{array}$
Then we obtain the next lemma.
Lemma 4.3.
The following elements $g_{1},\dots,g_{10}$ and their inverses $g_{1}^{-1},\dots,g_{10}^{-1}$ are contained in $PJ_{4}$.
$\begin{array}[]{ll}g_{1}=a_{1}s_{14}&g_{1}^{-1}=a_{16}s_{14}\\ g_{2}=a_{2}&g_{2}^{-1}=a_{11}\\ g_{3}=a_{3}s_{14}&g_{3}^{-1}=a_{14}s_{14}\\ g_{4}=a_{4}s_{14}&g_{4}^{-1}=a_{9}s_{14}\\ g_{5}=a_{5}&g_{5}^{-1}=a_{20}\\ g_{6}=a_{6}s_{14}&g_{6}^{-1}=a_{15}s_{14}\\ g_{7}=a_{7}s_{14}&g_{7}^{-1}=a_{18}s_{14}\\ g_{8}=a_{8}&g_{8}^{-1}=a_{13}\\ g_{9}=a_{10}s_{14}&g_{9}^{-1}=a_{19}s_{14}\\ g_{10}=a_{12}s_{14}&g_{10}^{-1}=a_{17}s_{14}\end{array}$
Furthermore, with respect to the action $\Gamma$ on $\mathcal{C}$ in Proposition 4.2, they are all of the elements of $PJ_{4}$ whose translation lengths are at most 4 and actually equal to $4$. Precisely, if $d(e,g\cdot e)\leq 4$ holds for $g$ in $PJ_{4}$, then $g=g_{i}^{\pm 1}$ and $d(e,g_{i}\cdot e)=4$ for some $1\leq i\leq 10$.
Proof.
We first confirm that $g_{1},...,g_{10}$ and their inverses $g_{1}^{-1},\dots,g_{10}^{-1}$ are in $PJ_{4}$. The elements $a_{1},\cdots,a_{8},a_{10},a_{12}$ correspond to the elements in $S_{4}$ via $\pi$ as follows.
$\pi(a_{1})=\pi(s_{13}s_{24}s_{12}s_{34})=(13)(24)(12)(34)=(14)(23)$
$\pi(a_{2})=\pi(s_{13}s_{24}s_{12}s_{34})=(13)(24)(13)(24)=e$
$\pi(a_{3})=\pi(s_{13}s_{34}s_{23}s_{12})=(13)(34)(23)(12)=(14)(23)$
$\pi(a_{4})=\pi(s_{13}s_{34}s_{13}s_{23})=(13)(34)(13)(23)=(14)(23)$
$\pi(a_{5})=\pi(s_{23}s_{12}s_{23}s_{13})=(23)(12)(23)(13)=e$
$\pi(a_{6})=\pi(s_{23}s_{12}s_{24}s_{12})=(23)(12)(24)(12)=(14)(23)$
$\pi(a_{7})=\pi(s_{23}s_{34}s_{13}s_{34})=(23)(34)(13)(34)=(14)(23)$
$\pi(a_{8})=\pi(s_{24}s_{34}s_{23}s_{34})=(24)(34)(23)(34)=e$
$\pi(a_{10})=\pi(s_{24}s_{12}s_{23}s_{34})=(24)(12)(23)(34)=(14)(23)$
$\pi(a_{12})=\pi(s_{24}s_{23}s_{13}s_{34})=(24)(23)(13)(34)=(14)(23)$
Thus, by definition, the elements $g_{1},...,g_{10}$ are in $PJ_{4}$.
Next we show their inverses $g_{1}^{-1},\dots,g_{10}^{-1}$ can be described as in the statement of the lemma.
For $a_{1},\dots,a_{20}$, the following are obtained by straightforward calculations.
$\begin{array}[]{llll}a_{1}^{-1}=a_{17}&a_{2}^{-1}=a_{11}&a_{3}^{-1}=a_{19}&a_{4}^{-1}=a_{4}\\ a_{5}^{-1}=a_{20}&a_{6}^{-1}=a_{18}&a_{7}^{-1}=a_{15}&a_{8}^{-1}=a_{13}\\ a_{9}^{-1}=a_{9}&a_{10}^{-1}=a_{14}&a_{12}^{-1}=a_{16}&\end{array}$
For example,
$\begin{array}[]{ll}a_{1}a_{17}=s_{13}s_{24}s_{12}s_{34}s_{34}s_{12}s_{24}s_{13}=e,\\ a_{17}a_{1}=s_{34}s_{12}s_{24}s_{13}s_{13}s_{24}s_{12}s_{34}=e.\end{array}$
We can also obtain
$\begin{array}[]{ll}s_{14}a_{1}=a_{12}s_{14}&s_{14}a_{2}=a_{11}s_{14}\\ s_{14}a_{3}=a_{10}s_{14}&s_{14}a_{4}=a_{9}s_{14}\\ s_{14}a_{5}=a_{8}s_{14}&s_{14}a_{6}=a_{7}s_{14}\\ s_{14}a_{7}=a_{6}s_{14}&s_{14}a_{8}=a_{5}s_{14}\\ s_{14}a_{9}=a_{4}s_{14}&s_{14}a_{10}=a_{3}s_{14}\\ s_{14}a_{11}=a_{2}s_{14}&s_{14}a_{12}=a_{1}s_{14}\end{array}$
by simple calculation in the same way as the following.
$s_{14}a_{1}$ $=s_{14}s_{13}s_{24}s_{12}s_{34}$
$=s_{24}s_{14}s_{24}s_{12}s_{34}$
$=s_{24}s_{13}s_{14}s_{12}s_{34}$
$=s_{24}s_{13}s_{34}s_{14}s_{34}$
$=s_{24}s_{13}s_{34}s_{12}s_{14}$
$=s_{24}s_{13}s_{12}s_{34}s_{14}$
$=s_{24}s_{23}s_{13}s_{34}s_{14}$
$=a_{12}s_{14}$
Thus, when we set $g_{1},\dots,g_{20}$ as in the statement, we have
$\begin{array}[]{ll}g_{1}^{-1}=a_{16}s_{14}&g_{2}^{-1}=a_{11}\\ g_{3}^{-1}=a_{14}s_{14}&g_{4}^{-1}=a_{9}s_{14}\\ g_{5}^{-1}=a_{20}&g_{6}^{-1}=a_{15}s_{14}\\ g_{7}^{-1}=a_{18}s_{14}&g_{8}^{-1}=a_{13}\\ g_{9}^{-1}=a_{19}s_{14}&g_{10}^{-1}=a_{17}s_{14}.\end{array}$
by
$g_{1}^{-1}$ $=(a_{1}s_{14})^{-1}=(s_{14}a_{12})^{-1}$
$=a_{12}^{-1}s_{14}=a_{16}s_{14}$
and so on.
Finally, with respect to the action $\Gamma$ on $\mathcal{C}$ in Proposition 4.2, we prove that $g_{1},\dots,g_{10}$ and their inverse are all of the elements of $PJ_{4}$ whose translation lengths are equal to $4$, i.e., if $d(e,g\cdot e)\leq 4$ holds for $g$ in $PJ_{4}$, then $g=g_{i}^{\pm 1}$ and $d(e,g_{i}\cdot e)=4$ for some $1\leq i\leq 10$.
Suppose that an element $g$ of $PJ_{4}$ has translation length at most $4$, i.e. $d(e,g\cdot e)\leq 4$ holds. This means that the element $g\cdot e$ in $J_{4}^{\prime}$ has word length at most 4.
All the elements of $J_{4}^{\prime}$ whose word lengths are $4$ or less can be listed up by brute force enumerations. See Tables 1 and 2. Actually, there are 40 elements of length 3 and 105 elements of length 4.
If $a:=g\cdot e$ in $J_{4}^{\prime}$, then, by the next claim, it lies on the image of $PJ_{4}$ under the map $\mathfrak{p}$, and so $\pi(a)\in\{e,(14)(23)\}$.
Claim 1.
Let $\pi$ denotes the natural surjection from $J_{4}$ to $S_{4}$, and $\mathfrak{p}$ the map defined by
$\mathfrak{p}:PJ_{4}$ $\longrightarrow J_{4}^{\prime}$
$g$$\longmapsto\begin{cases}g&g\in J_{4}^{\prime}\\ gs_{14}&g\notin J_{4}^{\prime}.\end{cases}$
Then, the following holds.
$\mathfrak{p}(PJ_{4})=\left(\pi|_{J_{4}^{\prime}}\right)^{-1}(\{e,(14)(23)\})$
Proof.
For any elements $x$ in $\left(\pi|_{J_{4}^{\prime}}\right)^{-1}(\{e,(14)(23)\})$, $\pi(x)=e$ or $(14)(23)$. Since $x$ lies on $J_{4}^{\prime}$, $x$ does not contain $s_{14}$ as a word. In the former case, $x$ is in $PJ_{4}$ by definition. Thus, $\mathfrak{p}(x)=x$. In the later case, the following equalities hold.
$\begin{array}[]{ll}\pi(x)&=(14)(23)\\ \pi(x)\pi(s_{14})&=(14)(23)(14)(23)\\ \pi(xs_{14})&=e\end{array}$
Therefore, $xs_{14}$ is in $PJ_{4}$ and $xs_{14}$ is not in $J_{4}^{\prime}$. It follows $\mathfrak{p}(xs_{14})=xs_{14}^{2}=x$. Conversely, each element $x$ in $\mathfrak{p}(PJ_{4})$, $x$ lies on $J_{4}^{\prime}$. By the construction of $\mathfrak{p}$, there exists $y$ in $PJ_{4}$ such that $x=y$ or $x=ys_{14}$. It follows that $\pi|_{J_{4}^{\prime}}(x)=e$ or $\pi|_{J_{4}^{\prime}}(x)=(14)(23)$. Therefore, $x$ is in $\left(\pi|_{J_{4}^{\prime}}\right)^{-1}(\{e,(14)(23)\})$. ∎
Among the elements in Tables 1 and 2, by direct checking whether the images of the elements under $\pi$ are $e$ or $(14)(23)$, we have only the elements $a_{1},\cdots,a_{20}$ given in the above. (Remark that $a_{16}=s_{34}s_{13}s_{23}s_{24}$ appears as $s_{34}s_{13}s_{24}s_{34}$ in Table 2.)
Consequently, all of elements of $PJ_{4}$ whose translation lengths are at most $4$ are shown to be $g_{1},...,g_{10}$ and their inverse elements $g_{1}^{-1},\dots,g_{10}^{-1}$. ∎
$\begin{array}[]{cccccc}s_{13}s_{23}s_{34}&s_{13}s_{23}s_{24}&s_{13}s_{24}s_{12}&s_{13}s_{24}s_{13}&s_{13}s_{24}s_{23}&s_{13}s_{34}s_{23}\\ s_{13}s_{34}s_{13}&s_{13}s_{34}s_{12}&s_{23}s_{13}s_{24}&s_{23}s_{13}s_{23}&s_{23}s_{12}s_{23}&s_{23}s_{12}s_{24}\\ s_{23}s_{12}s_{34}&s_{23}s_{34}s_{13}&s_{23}s_{34}s_{23}&s_{23}s_{34}s_{24}&s_{24}s_{34}s_{13}&s_{24}s_{34}s_{12}\\ s_{24}s_{12}s_{24}&s_{24}s_{12}s_{23}&s_{24}s_{12}s_{13}&s_{24}s_{13}s_{24}&s_{24}s_{13}s_{34}&s_{24}s_{13}s_{12}\\ s_{34}s_{24}s_{12}&s_{34}s_{24}s_{34}&s_{34}s_{23}s_{34}&s_{34}s_{23}s_{12}&s_{34}s_{23}s_{13}&s_{34}s_{13}s_{34}\\ s_{34}s_{13}s_{24}&s_{34}s_{13}s_{23}&s_{12}s_{34}s_{23}&s_{12}s_{34}s_{24}&s_{12}s_{24}s_{13}&s_{12}s_{24}s_{12}\\ s_{12}s_{24}s_{34}&s_{12}s_{23}s_{34}&s_{12}s_{23}s_{12}&s_{12}s_{23}s_{13}&&\end{array}$
Table 1. Elements of $J_{4}^{\prime}$ of length 3
$\begin{array}[]{cccc}s_{13}s_{23}s_{34}s_{12}&s_{13}s_{23}s_{34}s_{13}&s_{13}s_{23}s_{34}s_{23}&s_{13}s_{23}s_{34}s_{24}\\ s_{13}s_{24}s_{34}s_{13}&s_{13}s_{24}s_{34}s_{12}&s_{13}s_{24}s_{12}s_{24}&s_{13}s_{24}s_{12}s_{13}\\ s_{13}s_{24}s_{12}s_{13}&s_{13}s_{24}s_{13}s_{24}&s_{13}s_{24}s_{13}s_{34}&s_{13}s_{24}s_{13}s_{12}\\ s_{13}s_{34}s_{24}s_{12}&s_{13}s_{34}s_{24}s_{34}&s_{13}s_{34}s_{23}s_{34}&s_{13}s_{34}s_{23}s_{12}\\ s_{13}s_{34}s_{23}s_{13}&s_{13}s_{34}s_{13}s_{34}&s_{13}s_{34}s_{13}s_{24}&s_{13}s_{34}s_{13}s_{23}\\ s_{13}s_{12}s_{34}s_{23}&s_{13}s_{12}s_{34}s_{24}&s_{23}s_{13}s_{24}s_{13}&s_{23}s_{13}s_{24}s_{12}\\ s_{23}s_{13}s_{24}s_{34}&s_{23}s_{12}s_{13}s_{34}&s_{23}s_{12}s_{13}s_{12}&s_{23}s_{12}s_{23}s_{12}\\ s_{23}s_{12}s_{23}s_{34}&s_{23}s_{12}s_{23}s_{24}&s_{23}s_{12}s_{24}s_{12}&s_{23}s_{12}s_{24}s_{13}\\ s_{23}s_{12}s_{24}s_{23}&s_{23}s_{34}s_{12}s_{23}&s_{23}s_{34}s_{12}s_{13}&s_{23}s_{34}s_{13}s_{24}\\ s_{23}s_{34}s_{13}s_{34}&s_{23}s_{34}s_{13}s_{12}&s_{23}s_{34}s_{23}s_{12}&s_{23}s_{34}s_{23}s_{34}\\ s_{23}s_{34}s_{23}s_{24}&s_{23}s_{24}s_{23}s_{12}&s_{23}s_{24}s_{23}s_{13}&s_{24}s_{34}s_{13}s_{34}\\ s_{24}s_{34}s_{13}s_{24}&s_{24}s_{34}s_{13}s_{23}&s_{24}s_{12}s_{34}s_{23}&s_{24}s_{12}s_{34}s_{24}\\ s_{24}s_{12}s_{24}s_{13}&s_{24}s_{12}s_{24}s_{12}&s_{24}s_{12}s_{24}s_{34}&s_{24}s_{12}s_{23}s_{34}\\ s_{24}s_{12}s_{23}s_{12}&s_{24}s_{12}s_{23}s_{13}&s_{24}s_{13}s_{23}s_{34}&s_{24}s_{13}s_{23}s_{24}\\ s_{24}s_{13}s_{24}s_{12}&s_{24}s_{13}s_{24}s_{13}&s_{24}s_{13}s_{24}s_{23}&s_{24}s_{13}s_{34}s_{23}\\ s_{24}s_{13}s_{34}s_{13}&s_{24}s_{13}s_{34}s_{12}&s_{24}s_{23}s_{13}s_{24}&s_{24}s_{23}s_{13}s_{23}\\ s_{34}s_{24}s_{12}s_{23}&s_{34}s_{24}s_{12}s_{24}&s_{34}s_{24}s_{12}s_{34}&s_{34}s_{23}s_{24}s_{13}\\ s_{34}s_{23}s_{24}s_{23}&s_{34}s_{23}s_{34}s_{23}&s_{34}s_{23}s_{34}s_{13}&s_{34}s_{23}s_{34}s_{12}\\ s_{34}s_{23}s_{12}s_{24}&s_{34}s_{23}s_{12}s_{23}&s_{34}s_{23}s_{12}s_{13}&s_{34}s_{13}s_{12}s_{24}\\ s_{34}s_{13}s_{12}s_{34}&s_{34}s_{13}s_{34}s_{13}&s_{34}s_{13}s_{34}s_{23}&s_{34}s_{13}s_{34}s_{24}\\ s_{34}s_{13}s_{24}s_{13}&s_{34}s_{13}s_{24}s_{12}&s_{34}s_{13}s_{24}s_{34}&s_{34}s_{12}s_{13}s_{34}\\ s_{34}s_{12}s_{13}s_{12}&s_{12}s_{34}s_{23}s_{12}&s_{12}s_{34}s_{23}s_{34}&s_{12}s_{34}s_{23}s_{24}\\ s_{12}s_{24}s_{23}s_{12}&s_{12}s_{24}s_{23}s_{13}&s_{12}s_{24}s_{13}s_{34}&s_{12}s_{24}s_{13}s_{24}\\ s_{12}s_{24}s_{13}s_{23}&s_{12}s_{24}s_{12}s_{23}&s_{12}s_{24}s_{12}s_{24}&s_{12}s_{24}s_{12}s_{34}\\ s_{12}s_{23}s_{24}s_{13}&s_{12}s_{23}s_{24}s_{23}&s_{12}s_{23}s_{34}s_{23}&s_{12}s_{23}s_{34}s_{13}\\ s_{12}s_{23}s_{34}s_{12}&s_{12}s_{23}s_{12}s_{24}&s_{12}s_{23}s_{12}s_{23}&s_{12}s_{23}s_{12}s_{13}\\ s_{12}s_{13}s_{12}s_{24}&&&\end{array}$
Table 2. Elements of $J_{4}^{\prime}$ of length 4
Let $\widetilde{D}$ be the polygon on $\mathcal{C}$ defined by following, which is is visualized in Figure 4.
$\displaystyle\widetilde{D}:=\bigcap_{i=e}^{20}\left\{x\in\mathcal{C}\mid\ d(e,x)\leq d(x,a_{i})\right\}$
[…]
Figure 4.
From Figure 4, we see the following.
Lemma 4.4.
The word length of the vertices $\mathrm{int}\widetilde{D}$ are at most $3$. The vertices on $\partial\widetilde{D}$ whose word length is $3$ are only $s_{13}s_{24}s_{23}$, $s_{23}s_{34}s_{12}$, $s_{24}s_{13}s_{23}$, $s_{34}s_{13}s_{12}$, $s_{12}s_{24}s_{34}$.
We denote the five vertices above by $v_{1}$, …, $v_{5}$, respectively.
We remark that there are ten sides of $\widetilde{D}$ which are not edges in the Cayley complex $\mathcal{C}$. These sides are the diagonals of some 2-cells (quadrangles) in $\mathcal{C}$. The interior angles around $v_{1}$, …, $v_{5}$ are $\frac{2\pi}{5}$. The interior angles around $s_{13}s_{23}$, $s_{13}s_{12}$, $s_{24}s_{34}$, $s_{24}s_{23}$, $s_{34}s_{12}$ are $\frac{4\pi}{5}$. The interior angles around $s_{12}s_{23}$, $s_{13}s_{24}$, $s_{13}s_{34}$, $s_{23}s_{12}$, $s_{23}s_{34}$, $s_{24}s_{12}$, $s_{24}s_{13}$, $s_{34}s_{23}$, $s_{34}s_{13}$, $s_{12}s_{24}$ are $\frac{3\pi}{5}$.
Lemma 4.5.
The polygon $\widetilde{D}$ is the Dirichlet polygon with center $e$ for $PJ_{4}$ as a subgroup of $\mathrm{Isom}\;\mathcal{C}$.
Proof.
Since $PJ_{4}$ is nontrivial, we consider
$D(e)=\bigcap_{g\neq e}H_{g}(e).$
Here $H_{g}(e)$ denotes the half space defined by $\{x\in\mathcal{C}\mid d(x,e)<d(x,g\cdot e)\}$ for $g$ in $PJ_{4}$.
[…]
Let $g$ be a non-trivial element of $PJ_{4}$. By Lemma 4.3, if $g$ satisfies that $d(e,g\cdot e)$ is $4$ or less, then it is equal to $4$ and $g=g_{i}^{\pm 1}$ for some $i$. Therefore, $H_{g}(e)$ is equal to $H_{g_{i}^{\pm 1}}(e)$ for some $i$ when $d(e,g\cdot e)$ is 4 or less.
If $d(e,g\cdot e)$ is $6$ or more, we obtain the following.
$\{x\mid d(e,x)<3\}\subset H_{g}(e)$
[…]
It remains to show that $d(e,g\cdot e)\neq 5$ on $\mathcal{C}$ for any $g$ in $PJ_{4}$.
Recall that the image of the standard generators $s_{12}$, $s_{23}$, $s_{34}$, $s_{13}$, $s_{24}$ of $J_{4}^{\prime}$ under $\pi$ are the transpositions $(12)$, $(23)$, $(34)$, $(13)$, $(24)$ in $S_{4}$, and also that $\pi(s_{14})=(14)(23)$. See Section 2.
Let $g$ be an element in $PJ_{4}$. It means that $\pi(g)=e$ in $S_{4}$.
If $g$ is contained in $J_{4}^{\prime}$, then $g\cdot e=g$ is in $J_{4}^{\prime}$. Thus $d(e,g\cdot e)=d(e,g)=5$ implies $g$ has word length 5 with respect to the generators $s_{12}$, $s_{23}$, $s_{34}$, $s_{13}$, $s_{24}$ in $J_{4}^{\prime}$. It would imply that $\pi(g)$ is an odd permutation, contradicting that $\pi(g)=e$.
If $g$ is not contained in $J_{4}^{\prime}$, then $g\cdot e=gs_{14}$ is in $J_{4}^{\prime}$ by definition of the action $\Gamma$. Thus $d(e,g\cdot e)=d(e,gs_{14})=5$ implies $gs_{14}$ has word length 5 with respect to the generators $s_{12}$, $s_{23}$, $s_{34}$, $s_{13}$, $s_{24}$ in $J_{4}^{\prime}$, that is, $\pi(gs_{14})$ must be an odd permutation. Since $\pi(s_{14})=(14)(23)$, it would imply that $\pi(g)$ is an odd permutation, contradicting that $\pi(g)=e$.
As a result, $D(e)$ contains the interior of $\widetilde{D}$, and it concludes that the polygon $\widetilde{D}$ is the Dirichlet polygon with center $e$ for $PJ_{4}$ as a subgroup of $\mathrm{Isom}\;\mathcal{C}$. ∎
In the following, the side of $\partial\widetilde{D}$ with endpoints $v,v^{\prime}$ is denoted by $s(v,v^{\prime})$.
Proof of Theorem 1.1.
By Lemma 4.5, the interior of $\widetilde{D}$ is the Dirichlet domain for the action $\Gamma$ by $PJ_{4}$. Since the group which acts on a topological space freely and cocompactly is discontinuous, togethter with Proposition 4.2, $PJ_{4}$ is shown to be discontinuous. By Proposition 4.1, $\mathcal{C}$ is geodesically conected, geodesically complete, and finitely compact. Thus, applying Proposition 3.1 to the interior of $\widetilde{D}$, we obtain that $\widetilde{D}$ is a locally finite fundamental polygon for $PJ_{4}$. By Proposition 3.2, $PJ_{4}$ is generated by the elements $g_{1},\dots,g_{10}$ and their inverses.
Next, we will obtain the relations with respect to $g_{1},...,g_{10}$ by applying Proposition 3.3. The generators $g_{1},...,g_{10}$ of $PJ_{4}$ give identifications on (the sides of) the polygon $\widetilde{D}$. For example, by
$g_{1}(s_{34}s_{12})$ $=a_{1}s_{14}(s_{34}s_{12})s_{14}$
$=(s_{13}s_{24}s_{12}s_{34})s_{14}s_{14}(s_{12}s_{34})$
$=s_{13}s_{24}s_{12}s_{34}(s_{12}s_{34})$
$=s_{13}s_{24}s_{12}s_{34}(s_{34}s_{12})$
$=s_{13}s_{24},$
$g_{1}(s_{34}s_{13})$ $=a_{1}s_{14}(s_{34}s_{13})s_{14}$
$=(s_{13}s_{24}s_{12}s_{34})s_{14}s_{14}(s_{12}s_{24})$
$=s_{13}s_{24}s_{12}s_{34}(s_{12}s_{24})$
$=s_{13}s_{24}s_{12}(s_{12}s_{34}s_{24})$
$=s_{13}s_{24}(s_{34}s_{24})$
$=s_{13}s_{24}(s_{24}s_{23})$
$=s_{13}s_{23},$
we see that
$g_{1}:s(s_{34}s_{12},s_{34}s_{13})\mapsto s(s_{13}s_{24},s_{13}s_{23}).$
In the same way, we have the following.
$\begin{array}[]{rll}g_{2}&:s(s_{24}s_{13},s_{24}s_{13}s_{23})&\mapsto s(s_{13}s_{24},s_{13}s_{24}s_{23})\\ g_{3}&:s(s_{34}s_{23},s_{34}s_{23}s_{13})&\mapsto s(s_{13}s_{34},s_{13}s_{34}s_{24})\\ g_{4}&:s(s_{24}s_{34}),s_{24}s_{12})&\mapsto s(s_{13}s_{34},s_{13}s_{12})\\ g_{5}&:s(s_{13}s_{23},s_{12}s_{23})&\mapsto s(s_{23}s_{12},s_{23}s_{13})\\ g_{6}&:s(s_{34}s_{13},s_{34}s_{13}s_{12})&\mapsto s(s_{23}s_{12},s_{23}s_{12}s_{34})\\ g_{7}&:s(s_{12}s_{24},s_{12}s_{24}s_{34})&\mapsto s(s_{23}s_{34},s_{23}s_{34}s_{12})\\ g_{8}&:s(s_{24}s_{23},s_{34}s_{23})&\mapsto s(s_{23}s_{34},s_{24}s_{34})\\ g_{9}&:s(s_{12}s_{23},s_{12}s_{23}s_{24})&\mapsto s(s_{24}s_{12},s_{24}s_{12}s_{13})\\ g_{10}&:s(s_{12}s_{34},s_{12}s_{24})&\mapsto s(s_{24}s_{13},s_{24}s_{23})\end{array}$
It follows that $g_{1}$, …$g_{10}$ give identifications on $\widetilde{D}$. The identified polygon $\widetilde{D}^{*}$ under the identifications $g_{1}$, …, $g_{10}$ is a closed surface.
In the following, we will confirm that $\widetilde{D}$ is complete and satisfies the cycle condition, and so $\widetilde{D}$ is a Poincaré polygon.
Let $u_{1}$ be the vertex $s_{13}s_{24}s_{23}$ and $s_{1}$ the side $s(s_{13}s_{24}s_{23},s_{13}s_{24})$. Then $g_{2}^{-1}$ maps $u_{1}$ to the vertices $u_{2}:=s_{24}s_{13}s_{23}$ and $s_{1}$ to the side $s(s_{24}s_{13}s_{23},s_{24}s_{13})$. Let $s_{2}$ be the side $s(s_{24}s_{13}s_{23},s_{24}s_{12})$ which is the edge sharing $u_{2}$ with $g_{2}^{-1}(s_{1})$. Then, by $g_{9}^{-1}$, the vertex $u_{2}=s_{24}s_{13}s_{23}=s_{24}s_{12}s_{13}$ and the side $s_{2}$ are mapped to the vertex $u_{3}:=s_{12}s_{23}s_{24}=s_{12}s_{24}s_{34}$ and $s(s_{12}s_{23}s_{24},s_{12}s_{23})=s(s_{12}s_{24}s_{34},s_{12}s_{23})$. Let $s_{3}:=s(s_{12}s_{24}s_{34},s_{12}s_{24})$ be the side sharing $u_{3}$ with $g_{9}^{-1}(s_{2})$. The generator $g_{7}$ maps the vertex $u_{3}$ and the side $s_{3}$ to $u_{4}:=s_{23}s_{34}s_{12}$ and $s(s_{23}s_{34}s_{12},s_{23}s_{34})$. Let $s_{4}$ be the side $s(s_{23}s_{34}s_{12},s_{23}s_{12})=s(s_{23}s_{12}s_{34},s_{23}s_{12})$ which is the edge sharing $u_{4}$ with $g_{7}(s_{3})$. By the generator $g_{6}^{-1}$, the vertex $u_{4}$ and the side $s_{4}$ are mapped to $u_{5}:=s_{34}s_{13}s_{12}$ and $s(s_{34}s_{13}s_{12},s_{34}s_{13})$. Let $s_{5}$ be the side $s(s_{34}s_{13}s_{12},s_{34}s_{23})=s(s_{34}s_{23}s_{13},s_{34}s_{23})$, which is the side sharing $u_{5}$ with $g_{6}^{-1}(s_{4})$. The generator $g_{3}$ maps the vertex $u_{5}$ to $u_{1}$, and $s_{5}$ to the side $s(s_{13}s_{34}s_{24},s_{13}s_{34})=s(s_{13}s_{24}s_{23},s_{13}s_{34})$, which is the edge on $\partial\widetilde{D}$ sharing $u_{1}$ with $s_{1}$.
Therefore, the sequence of generators
$\{g_{2}^{-1},g_{9}^{-1},g_{7},g_{6}^{-1},g_{3}\}$
is a cycle of generators and the sequence of vertices $\{u_{1},\dots,u_{5}\}$ is a cycle of vertices. Moreover, all of the angles $\alpha(u_{i})$, made by the two sides around $u_{i}$ and measured from inside $\widetilde{D}$, are equal to $\frac{2\pi}{5}$. Thus, we obtain $\sum_{i=e}^{5}\alpha(u_{i})=2\pi$.
In the same way, we obtain the other cycles as follows.
$\{g_{4}^{-1},g_{8}^{-1},g_{3}\},\qquad\{s_{13}s_{34},s_{24}s_{34},s_{34}s_{23}\}$
The angles $\alpha(s_{13}s_{34})$ and $\alpha(s_{34}s_{23})$ are equal to $\frac{3\pi}{5}$ and the angle $\alpha(s_{24}s_{34})$ is $\frac{4\pi}{5}$, and so, the sum of the angles is equal to $2\pi$.
$\{g_{4}^{-1},g_{9}^{-1},g_{5}\},\qquad\{s_{13}s_{12},s_{24}s_{12},s_{12}s_{23}\}$
The angle $\alpha(s_{13}s_{12})$ is $\frac{4\pi}{5}$ and the angles $\alpha(s_{24}s_{12})$ and $\alpha(s_{12}s_{23})$ are equal to $\frac{3\pi}{5}$, and so, the sum of the angles is equal to $2\pi$.
$\{g_{6}^{-1},g_{1},g_{5}\},\qquad\{s_{23}s_{12},s_{34}s_{13},s_{13}s_{23}\}$
The angles $\alpha(s_{23}s_{12})$ and $\alpha(s_{34}s_{13})$ are equal to $\frac{3\pi}{5}$ and the angle $\alpha(s_{13}s_{23})$ is $\frac{4\pi}{5}$, and so, the sum of the angles is equal to $2\pi$.
$\{g_{7}^{-1},g_{10},g_{8}\},\qquad\{s_{23}s_{34},s_{12}s_{24},s_{24}s_{23}\}$
The angles $\alpha(s_{23}s_{34})$ and $\alpha(s_{12}s_{24})$ are equal to $\frac{3\pi}{5}$ and the angle $\alpha(s_{24}s_{23})$ is $\frac{4\pi}{5}$, and so, the sum of the angles is equal to $2\pi$.
$\{g_{2},g_{1}^{-1},g_{10}\},\qquad\{s_{24}s_{13},s_{13}s_{24},s_{34}s_{12}\}$
The angles $\alpha(s_{24}s_{13})$ and $\alpha(s_{13}s_{24})$ are equal to $\frac{3\pi}{5}$ and the angle $\alpha(s_{34}s_{12})$ is $\frac{4\pi}{5}$, and so, the sum of the angles is equal to $2\pi$.
It concludes that $\widetilde{D}$ satisfies the cycle condition, and so $\widetilde{D}$ is a Poincaré polygon. Applying Proposition 3.3, the compositions of generators $g_{3}g_{6}^{-1}g_{7}g_{9}^{-1}g_{2}^{-1}$, $g_{3}g_{8}^{-1}g_{4}^{-1}$, $g_{5}g_{9}^{-1}g_{4}^{-1}$, $g_{5}g_{1}g_{6}^{-1}$, $g_{8}g_{10}g_{7}^{-1}$, $g_{10}g_{1}^{-1}g_{2}$ form a complete set of relations. As a result, we obtain the presentation
$\left\langle g_{1},\cdots,g_{10}\left|\begin{array}[]{l}g_{3}g_{6}^{-1}g_{7}g_{9}^{-1}g_{2}^{-1}=g_{3}g_{8}^{-1}g_{4}^{-1}=g_{5}g_{9}^{-1}g_{4}^{-1}\\ =g_{5}g_{1}g_{6}^{-1}=g_{8}g_{10}g_{7}^{-1}=g_{10}g_{1}^{-1}g_{2}=e\end{array}\right.\right\rangle$
Moreover, by substituting $g_{1}=g_{2}g_{10}$, $g_{5}=g_{4}g_{9}$, $g_{6}=g_{5}g_{1}$, $g_{7}=g_{8}g_{10}$, $g_{3}=g_{4}g_{8}$ to $g_{3}g_{6}^{-1}g_{7}g_{9}^{-1}g_{2}^{-1}=e$, we obtain the following presentation.
$\langle g_{2},g_{4},g_{8},g_{9},g_{10}\mid g_{2}g_{9}g_{10}^{-1}g_{8}^{-1}g_{4}g_{9}g_{2}g_{10}g_{8}^{-1}g_{4}^{-1}=e\rangle$
We remark that, by calculating the Euler characteristic, we see that the surface obtained as $\widetilde{D}^{*}$ is homeomorphic to the connected sum of the five projective plane.
5. Proof of corollaries
In this section, we give proofs of corollaries.
Proof of Corollary 1.2.
Set
$\begin{array}[]{l}G:=\langle\alpha,\beta,\gamma,\delta,\epsilon\mid\alpha\gamma\epsilon\beta\epsilon\alpha^{-1}\delta^{-1}\beta\gamma\delta^{-1}=e\rangle,\\ G^{\prime}:=\langle g_{2},g_{4},g_{8},g_{9},g_{10}\mid g_{2}g_{9}g_{10}^{-1}g_{8}^{-1}g_{4}g_{9}g_{2}g_{10}g_{8}^{-1}g_{4}^{-1}=e\rangle.\end{array}$
Consider the following correspondence between the generators of $G$ and $G^{\prime}$.
$\begin{array}[]{ll}\alpha\longmapsto g_{4}g_{9}\\ \beta\longmapsto g_{2}g_{10}\\ \gamma\longmapsto g_{10}^{-1}\\ \delta\longmapsto g_{4}\\ \epsilon\longmapsto g_{8}^{-1}g_{4}g_{9}\end{array}$
Then the following calculation shows that the correspondence implies a well-defined homomorphism, say $f:G\to G^{\prime}$.
$\alpha\gamma\epsilon\beta\epsilon\alpha^{-1}\delta^{-1}\beta\gamma\delta^{-1}\longmapsto$ $g_{4}g_{9}g_{10}^{-1}g_{8}^{-1}g_{4}g_{9}g_{2}g_{10}g_{8}^{-1}g_{4}g_{9}(g_{4}g_{9})^{-1}g_{4}^{-1}g_{2}g_{10}g_{10}^{-1}g_{4}^{-1}$
$=g_{9}g_{10}^{-1}g_{8}^{-1}g_{4}g_{9}g_{2}g_{10}g_{8}^{-1}g_{4}^{-1}g_{2}=e$
This map $f$ is a surjection since the following hold.
$f(\beta\gamma)$ $=g_{2}g_{10}g_{10}^{-1}=g_{2}$
$f(\delta)$ $=g_{4}$
$f(\alpha\epsilon^{-1})$ $=g_{4}g_{9}g_{9}^{-1}g_{4}^{-1}g_{8}=g_{8}$
$f(\delta^{-1}\alpha)$ $=g_{4}^{-1}g_{4}g_{9}=g_{9}$
$f(\gamma^{-1})$ $=g_{10}$
In the same way, we can construct the inverse of $f$. The correspondence
$\begin{array}[]{ll}g_{2}\longmapsto\beta\gamma\\ g_{4}\longmapsto\delta\\ g_{8}\longmapsto\alpha\epsilon^{-1}\\ g_{9}\longmapsto\delta^{-1}\alpha\\ g_{10}\longmapsto\gamma^{-1}\end{array}$
induces a well-defined homomorphism $g:G^{\prime}\to G$ by the following.
$g_{10}^{-1}g_{8}^{-1}g_{4}g_{9}g_{2}g_{10}g_{8}^{-1}g_{4}^{-1}g_{2}g_{9}\longmapsto$ $\gamma\epsilon\alpha^{-1}\delta\delta^{-1}\alpha\beta\gamma\gamma^{-1}\epsilon\alpha^{-1}\delta^{-1}\beta\gamma\delta^{-1}\alpha$
$=\gamma\epsilon\beta\epsilon\alpha^{-1}\delta^{-1}\beta\gamma\delta^{-1}\alpha=e$
Then it follows that $g\circ f=id$ from the following.
$\begin{array}[]{ll}g\circ f(\alpha)&=g(g_{4}g_{9})=\delta\delta^{-1}\alpha=\alpha\\ g\circ f(\beta)&=g(g_{2}g_{10})=\beta\gamma\gamma^{-1}=\beta\\ g\circ f(\gamma)&=g(g_{10}^{-1})=(\gamma^{-1})^{-1}=\gamma\\ g\circ f(\delta)&=g(g_{4})=\delta\\ g\circ f(\epsilon)&=g(g_{8}^{-1}g_{4}g_{9})=\epsilon\alpha^{-1}\delta\delta^{-1}\alpha=\epsilon\end{array}$
As a result, $f$ is shown to be an isomorphism. ∎
Proof of Corollary 1.3.
Set
$\begin{array}[]{ll}G&:=\langle\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5}\mid\alpha_{1}^{2}\alpha_{2}^{2}\alpha_{3}^{2}\alpha_{4}^{2}\alpha_{5}^{2}=e\rangle,\\[12.0pt] G^{\prime}&:=\left\langle g_{1},\cdots,g_{10}\left|\begin{array}[]{l}g_{3}g_{6}^{-1}g_{7}g_{9}^{-1}g_{2}^{-1}=g_{3}g_{8}^{-1}g_{4}^{-1}=g_{5}g_{9}^{-1}g_{4}^{-1}\\ =g_{5}g_{1}g_{6}^{-1}=g_{8}g_{10}g_{7}^{-1}=g_{10}g_{1}^{-1}g_{2}=e\end{array}\right.\right\rangle\end{array}$
Consider the following correspondence.
$\begin{array}[]{ll}\alpha_{1}\longmapsto g_{1}^{-1}=g_{10}^{-1}g_{2}^{-1}\\ \alpha_{2}\longmapsto g_{2}g_{10}g_{5}^{-1}g_{8}g_{3}^{-1}=g_{2}g_{10}g_{9}^{-1}g_{4}^{-2}\\ \alpha_{3}\longmapsto g_{4}\\ \alpha_{4}\longmapsto g_{9}g_{10}^{-1}\\ \alpha_{5}\longmapsto g_{8}^{-1}g_{6}=g_{8}^{-1}g_{4}g_{9}g_{2}g_{10}\end{array}$
Then, by the next calculation, we have a well-defined homomorphism $f:G\to G^{\prime}$.
$\alpha_{1}^{2}\alpha_{2}^{2}\alpha_{3}^{2}\alpha_{4}^{2}\alpha_{5}^{2}\longmapsto$ $g_{1}^{-2}(g_{2}g_{10}g_{5}^{-1}g_{8}g_{3}^{-1})^{2}g_{4}^{2}(g_{9}g_{10}^{-1})^{2}(g_{8}^{-1}g_{6})^{2}$
$=g_{1}^{-2}(g_{1}g_{5}^{-1}g_{4}^{-1})^{2}g_{4}^{2}(g_{9}g_{10}^{-1})^{2}(g_{8}^{-1}g_{6})^{2}$
$=g_{1}^{-1}g_{5}^{-1}g_{4}^{-1}g_{1}(g_{5}^{-1}g_{4})g_{9}g_{10}^{-1}g_{9}(g_{10}^{-1}g_{8}^{-1})g_{6}g_{8}^{-1}g_{6}$
$=g_{1}^{-1}g_{5}^{-1}g_{4}^{-1}g_{1}((g_{9}^{-1})g_{9})g_{10}^{-1}g_{9}(g_{7}^{-1})g_{6}g_{8}^{-1}g_{6}$
$=(g_{1}^{-1}g_{5}^{-1})g_{4}^{-1}(g_{1}g_{10}^{-1})g_{9}g_{7}^{-1}g_{6}g_{8}^{-1}g_{6}$
$=g_{6}^{-1}g_{4}^{-1}g_{3}(g_{6}^{-1}g_{6})g_{8}^{-1}g_{6}$
$=g_{6}^{-1}g_{4}^{-1}g_{4}g_{8}g_{8}^{-1}g_{6}=e$
This map $f$ is a surjection from the following.
$f(\alpha_{2}\alpha_{3}^{2}\alpha_{4})$ $=g_{2}g_{10}g_{9}^{-1}g_{4}^{-2}g_{4}^{2}g_{9}g_{10}^{-1}=g_{2}$
$f(\alpha_{3})$ $=g_{4}$
$f(\alpha_{3}^{-1}\alpha_{2}^{-1}\alpha_{1}^{-2}\alpha_{5}^{-1})$ $=g_{4}^{-1}g_{4}^{2}g_{9}g_{10}^{-1}g_{2}^{-1}(g_{2}g_{10})^{2}g_{10}^{-1}g_{2}^{-1}g_{9}^{-1}g_{4}^{-1}g_{8}=g_{8}$
$f(\alpha_{3}^{-2}\alpha_{2}^{-1}\alpha_{1}^{-1})$ $=g_{4}^{-2}g_{4}^{2}g_{9}g_{10}^{-1}g_{2}^{-1}g_{2}g_{10}=g_{9}$
$f((\alpha_{2}\alpha_{3}^{2}\alpha_{4})^{-1}\alpha_{1}^{-1})$ $=g_{2}^{-1}g_{2}g_{10}=g_{10}$
On the other hand, the correspondence
$\begin{array}[]{lll}g_{2}&\longmapsto\alpha_{2}\alpha_{3}^{2}\alpha_{4}\\ g_{4}&\longmapsto\alpha_{3}\\ g_{8}&\longmapsto\alpha_{3}^{-1}\alpha_{2}^{-1}\alpha_{1}^{-2}\alpha_{5}^{-1}\\ g_{9}&\longmapsto\alpha_{3}^{-2}\alpha_{2}^{-1}\alpha_{1}^{-1}\\ g_{10}&\longmapsto\alpha_{4}^{-1}\alpha_{3}^{-2}\alpha_{2}^{-1}\alpha_{1}^{-1}\end{array}$
induces a well-defined homomorphism $g:G^{\prime}\to G$ by the following.
$\begin{array}[]{ll}&g_{2}g_{9}g_{10}^{-1}g_{8}^{-1}g_{4}g_{9}g_{2}g_{10}g_{8}^{-1}g_{4}^{-1}\\ &\qquad\qquad\longmapsto(\alpha_{2}\alpha_{3}^{2}\alpha_{4})(\alpha_{3}^{-2}\alpha_{2}^{-1}\alpha_{1}^{-1})(\alpha_{1}\alpha_{2}\alpha_{3}^{2}\alpha_{4})(\alpha_{5}\alpha_{1}^{2}\alpha_{2}\alpha_{3})\alpha_{3}(\alpha_{3}^{-2}\alpha_{2}^{-1}\alpha_{1}^{-1})\\ &\qquad\qquad\qquad(\alpha_{2}\alpha_{3}^{2}\alpha_{4})(\alpha_{4}^{-1}\alpha_{3}^{-2})\alpha_{2}^{-1}\alpha_{1}^{-1}(\alpha_{5}\alpha_{1}^{2}\alpha_{2}\alpha_{3})\alpha_{3}^{-1}\\ &\qquad\qquad\qquad=\alpha_{2}\alpha_{3}^{2}\alpha_{4}\alpha_{4}\alpha_{5}\alpha_{1}^{2}(\alpha_{2}\alpha_{3}\alpha_{3}\alpha_{3}^{-2}\alpha_{2}^{-1})\alpha_{1}^{-1}(\alpha_{2}\alpha_{2}^{-1})\alpha_{1}^{-1}\alpha_{5}\alpha_{1}^{2}\alpha_{2}\\ &\qquad\qquad\qquad=\alpha_{2}\alpha_{3}^{2}\alpha_{4}\alpha_{4}\alpha_{5}(\alpha_{1}^{2}\alpha_{1}^{-1}\alpha_{1}^{-1})\alpha_{5}\alpha_{1}^{2}\alpha_{2}\\ &\qquad\qquad\qquad=\alpha_{2}\alpha_{3}^{2}\alpha_{4}\alpha_{4}\alpha_{5}\alpha_{5}\alpha_{1}^{2}\alpha_{2}=e\end{array}$
Then it follows that $g\circ f=id$ as;
$\begin{array}[]{ll}g\circ f(\alpha_{1})=g(g_{10}^{-1}g_{2}^{-1})=\alpha_{1}(\alpha_{2}\alpha_{3}^{2}\alpha_{4}\alpha_{4}^{-1}\alpha_{3}^{-2}\alpha_{2}^{-1})=\alpha_{1}\\ g\circ f(\alpha_{2})=g(g_{2}g_{10}g_{9}^{-1}g_{4}^{-2})=\alpha_{2}(\alpha_{3}^{2}\alpha_{4}\alpha_{4}^{-1}\alpha_{3}^{-2})(\alpha_{2}^{-1}\alpha_{1}^{-1}\alpha_{1}\alpha_{2})(\alpha_{3}^{2}\alpha_{3}^{-2})=\alpha_{2}\\ g\circ f(\alpha_{3})=g(g_{4})=\alpha_{3}\\ g\circ f(\alpha_{4})=g(g_{9}g_{10}^{-1})=\alpha_{3}^{-2}\alpha_{2}^{-1}\alpha_{1}^{-1}\alpha_{1}\alpha_{2}\alpha_{3}^{2}=\alpha_{4}\\ g\circ f(\alpha_{5})=g(g_{8}^{-1}g_{4}g_{9}g_{2}g_{10})=\alpha_{5}(\alpha_{1}^{2}\alpha_{2}\alpha_{3}\alpha_{3}\alpha_{3}^{-2}\alpha_{2}^{-1}\alpha_{1}^{-1})\alpha_{2}(\alpha_{3}^{2}\alpha_{4}\alpha_{4}^{-1}\alpha_{3}^{-2})\alpha_{2}^{-1}\alpha_{1}^{-1}\\ \qquad\qquad\,=\alpha_{5}(\alpha_{1}\alpha_{2}\alpha_{2}^{-1}\alpha_{1}^{-1})=\alpha_{5}\end{array}$
This concludes that $f$ gives an isomorphism. ∎
Remark 2.
By virtue of geometric descriptions, we could find the correspondence between $\alpha_{i}$ and $g_{j}$ in the proof above as follows. The painted regions in Figure 5 give five Möbius bands embedded in $\mathcal{C}/PJ_{4}$. After removing them, we can rebuild the remaining parts to obtain a sphere with five holes. See Figure 6. Then we can confirm that $\mathcal{C}/PJ_{4}$ is homeomorphic to the connected sum #$\#_{5}\mathbb{RP}^{2}$ of five real projective planes. Setting the closed curve $\alpha_{2}$ in #$\#_{5}\mathbb{RP}^{2}$ shown in Figure 6, we obtain a correspondence $\alpha_{2}\longmapsto g_{8}^{-1}g_{4}g_{9}g_{2}g_{10}\in PJ_{4}$ by considering a lift of $\alpha_{2}$ in $\mathcal{C}$. In the same way, we can obtain the corresponding elements in $PJ_{4}$ for $\alpha_{1}$, $\alpha_{3}$, $\alpha_{4}$, $\alpha_{5}$.
[…]
Figure 5.
[…]
Figure 6.
Acknowledgments
We would like to thank Takuya Sakasai for his valuable feedback and Carl-Fredrik Nyberg Brodda for inspiring future research themes.

### v2

@@
 The kernel of this surjection is called the pure cactus group $PJ_{n}$ of degree $n$.
-In [9, Theorem 9], Henriques and Kamnitzer showed that $PJ_{n}$ is isomorphic to the fundamental group of the Deligne–Mumford compactification $\overline{M_{0,n+1}}(\mathbb{R})$ of the moduli space of real genus-zero curves with $n+1$ marked points.
+In [10, Theorem 9], Henriques and Kamnitzer showed that $PJ_{n}$ is isomorphic to the fundamental group of the Deligne–Mumford compactification $\overline{M_{0,n+1}}(\mathbb{R})$ of the moduli space of real genus-zero curves with $n+1$ marked points.
 In the degree four case, it is known that $\overline{M_{0,5}}(\mathbb{R})$ is homeomorphic to the connected sum of five real projective planes (see [6, Example 2.5]).
@@
 In this section, we give definitions of the cactus group and the pure cactus group algebraically.
-For a given integer $n\geq 2$, the cactus group of degree $n$, denoted by $J_{n}$, is defined as the group admitting a presentation with the generators $s_{p,q}$ for $1\leq p<q\leq n$ and the following relations.
+For any integer $n\geq 2$, we define the cactus group of degree $n$, denoted by $J_{n}$, as the group given by the presentation with generators $s_{p,q}$ for $1\leq p<q\leq n$, subject to the following relations.
   • $s_{p,q}^{2}=e$ for every $1\leq p<q\leq n$,
@@
   • $s_{p,q}s_{m,r}=s_{p+q-r,p+q-m}s_{p,q}$ for all $1\leq p<q\leq n$ and $1\leq m<r\leq n$ satisfying $[m,r]\subset[p,q]$.
-Here, $e$ denotes the trivial element of the group and $[p,q]$ denotes the set $\{p,p+1,\dots,q-1,q\}$ of integers for positive integers $p$, $q$ with $p<q$.
-Throughout the sequel, we denote the generators such as $s_{23}$ (without comma) in stead of $s_{2,3}$ for brevity.
+Here, $e$ denotes the identity element, and $[p,q]$ denotes the set ${p,p+1,\dots,q-1,q}$ of integers for positive integers $p$ and $q$ with $p<q$.
+Throughout the sequel, we write generators such as $s_{23}$ (without a comma) instead of $s_{2,3}$ for brevity.
 For the degree four case, the cuctus group $J_{4}$ has the following presentation.
 $\left\langle s_{12},s_{23},s_{34},s_{13},s_{24},s_{14}\left|\begin{array}[]{l}s_{12}^{2}=s_{23}^{2}=s_{34}^{2}=s_{13}^{2}=s_{24}^{2}=s_{14}^{2}=e,\\ s_{12}s_{34}=s_{34}s_{12},s_{12}s_{13}=s_{13}s_{23},s_{23}s_{24}=s_{24}s_{34},\\ s_{12}s_{14}=s_{14}s_{34},s_{23}s_{14}=s_{14}s_{23},s_{13}s_{14}=s_{14}s_{24}\end{array}\right.\right\rangle.$
-Similar to the braid group, elements of the cactus group can be represented by diagrams of vertical strands on the plane. Some of such diagrams for $J_{4}$ are exhibited in Figure 1.
+Similar to the braid group, elements of the cactus group can be represented by diagrams of vertical strands on the plane. Some such diagrams for $J_{4}$ are shown in Figure 1.
 $s_{14}$$s_{13}$$s_{12}$$s_{23}$$s_{34}$
@@
 Figure 1. Diagrams for some elements of $J_{4}$
-Due to this diagramatic expression, we see that the cactus group $J_{n}$ admits a natural projection $\pi:J_{n}\to S_{n}$ onto the symmetric group $S_{n}$ of degree $n$, and its kernel is called the pure cactus group of degree $n$, denoted by $PJ_{n}$.
-See [9, Subsection 3.1] or [7, Section 1] for more details.
-We also introduce certain subgroup used in the sections later.
-For each integer $n\geq 2$ and subset $S\subset[2,n]$, let $J_{n}^{S}$ be the subgroup of $J_{n}$ generated by the elements $s_{p,q}$ for $1\leq p<q\leq n$ and $q-p+1\in S$ and defined by the following relations.
+Due to this diagrammatic expression, we see that the cactus group $J_{n}$ admits a natural projection $\pi:J_{n}\to S_{n}$ onto the symmetric group $S_{n}$ of degree $n$, and its kernel is called the pure cactus group of degree $n$, denoted by $PJ_{n}$.
+See [10, Subsection 3.1] or [7, Section 1] for more details.
+We also introduce a certain subgroup that will be used in later sections.
+For each integer $n\geq 2$ and subset $S\subset[2,n]$, let $J_{n}^{S}$ be the subgroup of $J_{n}$ generated by the elements $s_{p,q}$ for $1\leq p<q\leq n$ with $q-p+1\in S$, and defined by the following relations.
   • $s_{p,q}^{2}=e$ for every $1\leq p<q\leq n$ satisfying $q-p+1\in S$,
@@
 In later sections, we mainly consider the group $J_{4}^{[2,3]}$, which we denote by $J_{4}^{\prime}$ in the rest of paper.
-It is a subgroup of $J_{4}$ generated by $s_{12},s_{23},s_{34},s_{13},s_{24}$ (except for $s_{14}$), and it has the following presentation.
+It is a subgroup of $J_{4}$ generated by $s_{12},s_{23},s_{34},s_{13},s_{24}$ (excluding $s_{14}$), and it has the following presentation.
 $\left\langle s_{12},s_{23},s_{34},s_{13},s_{24}\left|\begin{array}[]{l}s_{12}^{2}=s_{23}^{2}=s_{34}^{2}=s_{13}^{2}=s_{24}^{2}=e,\\ s_{12}s_{34}=s_{34}s_{12},s_{12}s_{13}=s_{13}s_{23},s_{23}s_{24}=s_{24}s_{34}\end{array}\right.\right\rangle.$
-We denote by $\mathcal{C}$ the Cayley complex of the group $J_{4}^{\prime}$ with respect to the presentation above.
+We denote by $\mathcal{G}$ the Cayley graph of the group $J_{4}^{\prime}$, and by $\mathcal{C}$ the Cayley complex of $J_{4}^{\prime}$ with respect to the presentation above.
+See, for instance, [9] for the definitions of these objects.
 3. Preliminary
-In this section, we prepare some properties of group action for the proof of the Theorem 1.1.
-The following two propositions are given in [14]. See [14] for definitions of the terms used.
-Proposition 3.1 ([14, Theorem 6.6.13]).
+In this section, we prepare some properties of group actions for the proof of Theorem 1.1.
+The following two propositions are given in [15]. See [15] for the definitions of the terms used.
+Proposition 3.1 ([15, Theorem 6.6.13]).
 Let $D(a)$ be the Dirichlet domain with center $a$ for a discontinuous group $G$ of isometries of a metric space $X$ such that
@@
 Then $D(a)$ is a locally finite fundamental domain for $G$.
-Proposition 3.2 ([14, Exercise 6.6, 2.]).
+Proposition 3.2 ([15, Exercise 6.6, 2.]).
 Let G be a group of isometries of a connected metric space $X$ with locally finite fundamental domain $R$. Then $G$ is generated by $\{g\in G\mid\overline{R}\cap g\overline{R}\neq\emptyset\}$.
-Next we recall the following well-known theorem originated in [13]. Our terminologies follows from [11].
-Proposition 3.3 (Poincaré’s polygon theorem, c.f. [11, p223]).
+Next, we recall the following well-known theorem, originally from [14]. Our terminology follows that of [12].
+Proposition 3.3 (Poincaré’s polygon theorem, c.f. [12, p223]).
 Let $D$ be a Poincaré polygon.
@@
 Then $G$ is discontinuous, $D$ is a fundamental polygon for $G$, and the cycle relations form the complete set of relations for $G$.
-In the following, we give explanations for the terms used in the theorem above.
-Let $D$ be a polygon on the hyperbolic plane $\mathbb{H}^{2}$.
-An identification on $D$ is a map which assigns, to each side $s$, a side $s^{\prime}$ and an isometry $A(s,s^{\prime})$ of $\mathbb{H}^{2}$ so that
+In the following, we explain the terms used in the theorem above.
+Let $D$ be a polygon in the hyperbolic plane $\mathbb{H}^{2}$.
+An identification on $D$ is a map that assigns, to each side $s$, a side $s^{\prime}$ and an isometry $A(s,s^{\prime})$ of $\mathbb{H}^{2}$ such that
   1. (i)
@@
   4. (iv)
-    for each side $s$, there exists a neighborhood $V$ of $s$ so that $A(s,s^{\prime})(V\cap D)\cap D=\emptyset$.
+    for each side $s$, there exists a neighborhood $V$ of $s$ such that $A(s,s^{\prime})(V\cap D)\cap D=\emptyset$.
 We call the isometries $A(s,s^{\prime})$ which identify the sides of the polygon $D$ identifying generators, or generators.
@@
 For two elements $x$, $x^{\prime}$ in $D^{*}$, we set a metric on $D^{*}$ as
-$d_{D^{*}}^{*}(x,x^{\prime}):=\inf\sum_{i=e}^{n}d_{\mathbb{H}^{2}}(z_{i},z_{i}^{\prime})$
+$d_{D^{*}}^{*}(x,x^{\prime}):=\inf\sum_{i=1}^{n}d_{\mathbb{H}^{2}}(z_{i},z_{i}^{\prime})$
 where the infimum is taken over all $n$ and over all $2n$-tuples of points of $\overline{D}$ such that $p(z_{1})=x$, $p(z_{i}^{\prime})=p(z_{i+1})$, and $p(z_{n}^{\prime})=x^{\prime}$.
@@
 Let $D$ be a complete polygon with identification.
-Let $z_{1}$ be some vertex of $D$.
-There are precisely two sides of $D$ which meet at $z_{1}$; we choose one of these and call it $s_{1}$.
+Let $z_{1}$ be a vertex of $D$.
+There are precisely two sides of $D$ meeting at $z_{1}$; we choose one of these and call it $s_{1}$.
 There is then a corresponding side $s_{1}^{\prime}$ and a generator $A_{1}=A(s_{1},s_{1}^{\prime})$.
-Set $z_{2}=A_{1}(z_{1})$ and observe that there is a unique other side $s_{2}$ which has $z_{2}$ as an end point.
+Set $z_{2}=A_{1}(z_{1})$ and observe that there is a unique other side $s_{2}$ having $z_{2}$ as an endpoint.
 There is a corresponding side $s_{2}^{\prime}$ and a generator $A_{2}=A(s_{2},s_{2}^{\prime})$.
-Set $z_{3}=A_{2}(z_{2})$ and observe that there is a unique other side $s_{3}$ which has $z_{3}$ as an end point.
-In this manner, we get a sequence $\{z_{i}\}$ of vertices, and a sequence $\{A_{i}\}$ of generators.
-Then there exists periodic sequence of generators $\{A_{1},\cdots A_{n}\}$ such that $A_{n}(\cdots A_{2}(A_{1}(z_{1})))=z_{1}$ which is called a cycle of generators.
-For a cycle of generators $\{A_{1},\dots,A_{n}\}$, there exists a sequence of vertices $\{z_{1},\dots,z_{n}\}$ in $D$ such that $z_{2}=A_{1}(z_{1})$, …, $z_{n}=A_{n-1}(z_{n-1})$, and $A_{n}(z_{n})=z_{1}$.
-The sequence of the vertices is called a cycle of vertices.
-Then, the polygon $D$ is said to satisfy the cycle condition if for each cycle of vertices $\{z_{1},\dots,z_{n}\}$, there exists an integer $\nu$ so that
+Set $z_{3}=A_{2}(z_{2})$ and observe that there is a unique other side $s_{3}$ having $z_{3}$ as an endpoint.
+In this way, we obtain a sequence $\{z_{i}\}$ of vertices, and a sequence $\{A_{i}\}$ of generators.
+Then there exists a periodic sequence of generators $\{A_{1},\dots,A_{n}\}$ such that
+$A_{n}(\cdots A_{2}(A_{1}(z_{1})))=z_{1}$
+called a cycle of generators. For a cycle of generators $\{A_{1},\dots,A_{n}\}$, there exists a sequence of vertices $\{z_{1},\dots,z_{n}\}$ in $D$ such that
+$z_{2}=A_{1}(z_{1}),\ \dots,\ z_{n}=A_{n-1}(z_{n-1}),\text{ and }A_{n}(z_{n})=z_{1}.$
+This sequence is called a cycle of vertices. Then, the polygon $D$ is said to satisfy the cycle condition if for each cycle of vertices $\{z_{1},\dots,z_{n}\}$, there exists an integer $\nu$ such that
   1. (vii)
-    $\nu\sum_{i=e}^{m}\alpha(z_{i})=2\pi$,
-where $\alpha(z_{i})$ is the angle around $x_{i}$ measured from inside $D$.
-Let $D$ be a polygon with identifications. The polygon $D$ is a Poincaré polygon if $D$ is complete and satisfies the cycle condition.
-For a cycle of generators $\{A_{1},\dots,A_{n}\}$, by Condition (vii) , the $\nu$ times of the composition $A_{n}\circ\cdots\circ A_{1}$ is the identity on $D$, i.e., we get a relation $(A_{n}\circ\cdots\circ A_{1})^{\nu}=e$ for that cycle.
-We call such relations the cycle relation.
-4. Proof of Theorem 1.1
-Before proving Theorem 1.1, we prepare the following key proposition.
+    $\nu\sum_{i=1}^{n}\alpha(z_{i})=2\pi$,
+where $\alpha(z_{i})$ is the interior angle at $z_{i}$ measured inside $D$.
+Let $D$ be a polygon with identifications. The polygon $D$ is called a Poincaré polygon if it is complete and satisfies the cycle condition.
+For a cycle of generators $\{A_{1},\dots,A_{n}\}$, by Condition (vii), the $\nu$-fold composition
+$(A_{n}\circ\cdots\circ A_{1})^{\nu}$
+is the identity on $D$, i.e., we have the relation
+$(A_{n}\circ\cdots\circ A_{1})^{\nu}=e$
+for that cycle. We call such relations the cycle relations.
+4. Key propositions
+In this section, we present two key propositions, Proposition 4.1 and Proposition 4.2, which are essential for the proof of Theorem 1.1.
 Proposition 4.1.
@@
 Proof.
-As shown in [7, Proposition 2.6], every element of the cactus group has a unique diagram of the normal form.
-This enable us to enumerate all the elements of $J_{4}^{\prime}$ with bounded word length, equivalently, to determine the local structure, vertices and edges, in $\mathcal{C}$ inside the bounded ball centered at $e$.
+As shown in [7, Proposition 2.6], every element of the cactus group has a unique normal form diagram.
+This enables us to enumerate all the elements of $J_{4}^{\prime}$ with bounded word length, equivalently, to determine the local structure, vertices and edges, in $\mathcal{C}$ inside the bounded ball centered at $e$.
 Note that the vertices of $\mathcal{C}$ are identified with elements in $J_{4}^{\prime}$.
-The vertices in $\mathcal{C}$ whose word length is $1$ are just generators;
+The vertices in $\mathcal{C}$ whose word length is $1$ are simply the generators;
 $\begin{array}[]{ccccc}s_{12}&s_{23}&s_{34}&s_{13}&s_{24}.\end{array}$
-The following are the vertices whose word length with respect to the generators above is $2$.
+The following vertices have word length 2 with respect to the generators above.
 $\begin{array}[]{ccccc}s_{12}s_{23}&s_{13}s_{23}&s_{13}s_{24}&s_{13}s_{34}&s_{13}s_{12}\\ s_{23}s_{12}&s_{23}s_{34}&s_{24}s_{34}&s_{24}s_{12}&s_{24}s_{13}\\ s_{24}s_{23}&s_{34}s_{23}&s_{34}s_{13}&s_{34}s_{12}&s_{12}s_{24}\end{array}$
-As shown Figure 2, There are $25$ edges ($1$-cells) connecting the above vertices (00-cells) in $\mathcal{C}$, and the $2$-cells (quadrangles) containing $e$ on $\mathcal{C}$ are
+As shown in Figure 2, there are 25 edges (1-cells) connecting the above vertices (0-cells) in $\mathcal{C}$, and the 2-cells (quadrangles) containing $e$ in $\mathcal{C}$ are
 $\begin{array}[]{l}\langle e,s_{12},s_{12}s_{13},s_{13}\rangle,\langle e,s_{13},s_{13}s_{12},s_{23}\rangle,\langle e,s_{23},s_{23}s_{24},s_{24}\rangle,\\ \langle e,s_{24},s_{24}s_{23},s_{34}\rangle,\langle e,s_{34},s_{34}s_{12},s_{12}\rangle.\end{array}$
@@
 Figure 2. Neighborhood of $e$ in the Cayley complex $\mathcal{C}$
-Therefore, $\mathcal{C}$ is a cell complex such that there are five quadrilateral $2$-cells surrounding each vertices.
-We consider the map $f$ corresponding each closed $2$-cell on $\mathcal{C}$ to the square (regular quadrilateral) in $\mathbb{H}^{2}$ whose internal angles are equal to $\frac{2\pi}{5}$.
+Therefore, $\mathcal{C}$ is a cell complex such that there are five quadrilateral $2$-cells surrounding each vertex.
+We consider the map $f$ corresponding to each closed $2$-cell on $\mathcal{C}$ to the square (regular quadrilateral) in $\mathbb{H}^{2}$ whose internal angles are equal to $\frac{2\pi}{5}$.
 The image of the map $f$ gives a uniform tiling (regular tessellation) of $\mathbb{H}^{2}$ of type $\{4,5\}$.
@@
 Figure 3. The $\{4,5\}$-tesselation of the hyperbolic plane. ([2, Figure 7])
-We see that this map is an isometry up to scaling $d(f(x),f(y))=Rd(x,y)$ with $R=\cosh^{-1}(\cot^{2}{\frac{\pi}{5}})$ by elementary calculation based on the hyperbolic trigonometry. ∎
-In other words, if we set a metric on $\mathcal{C}$ where the length of each edge is the constant $R$, then $\mathcal{C}$ is truly isometric to $\mathbb{H}^{2}$.
-Therefore, we can apply results for the hyperbolic plane to the Cayley complex $\mathcal{C}$.
-Now we consider an action of $PJ_{4}$ on $\mathcal{C}$. The following proposition was essentially obtained in [7].
+We see that this map is an isometry up to scaling, satisfying $d(f(x),f(y))=R\,d(x,y)$ with $R=\cosh^{-1}\left(\cot^{2}{\frac{\pi}{5}}\right)$ by an elementary calculation based on hyperbolic trigonometry.
+∎
+In other words, if we set a metric on $\mathcal{C}$ by assigning each edge a constant length $R$, then $\mathcal{C}$ is truly isometric to $\mathbb{H}^{2}$.
+Therefore, we can apply results for the hyperbolic plane directly to the Cayley complex $\mathcal{C}$.
+Next, we consider an action of $PJ_{4}$ on the Cayley complex $\mathcal{C}$. The following proposition was essentially obtained in [7].
 Proposition 4.2 ([7, Proof of Corollary 7.3]).
-The map $\Gamma$ induced from $\Gamma_{0}$ defined by the following implies an isometric action of $PJ_{4}$ on $\mathcal{C}$.
-Note that $\left(\mathcal{C}\right)^{(0)}$ is identified with $J_{4}^{\prime}$.
-$\Gamma_{0}:PJ_{4}\times\left(\mathcal{C}\right)^{(0)}$ $\longrightarrow\left(\mathcal{C}\right)^{(0)}$
+Let $\Gamma_{0}$ be the action of $PJ_{4}$ on $J_{4}^{\prime}$ defined as follows.
+$\Gamma_{0}:PJ_{4}\times J_{4}^{\prime}$ $\longrightarrow J_{4}^{\prime}$
 $(g,h)$$\longmapsto\begin{cases}gh&gh\in J_{4}^{\prime}\\ ghs_{14}&gh\notin J_{4}^{\prime}\end{cases}$
-Moreover, the action $\Gamma$ on $\mathcal{C}$ is free and cocompact.
-In the following, for simplicity, we identify $PJ_{4}$ and its image under the action $\Gamma$, and write $g\cdot x$ in stead of $\Gamma(g,x)$ for $g\in PJ_{4}$ and $x\in\mathcal{C}$.
+Then $\Gamma_{0}$ induces the maps $\Gamma_{1}$ and $\Gamma$, which give rise to isometric actions of $PJ_{4}$ on $\mathcal{G}$ and $\mathcal{C}$, respectively.
+Note that $\left(\mathcal{C}\right)^{(0)}=J_{4}^{\prime}$ is identified with $J_{4}^{\prime}$.
+Moreover, the actions $\Gamma_{1}$ and $\Gamma$ on $\mathcal{G}$ and $\mathcal{C}$ are free and cocompact.
+In the following, for simplicity, we identify $PJ_{4}$ with its image under the action $\Gamma$, and write $g\cdot x$ instead of $\Gamma(g,x)$ for $g\in PJ_{4}$ and $x\in\mathcal{C}$.
 In this setting, $PJ_{4}$ is regarded as a subgroup of $\mathrm{Isom}\;\mathcal{C}$.
-Now, let $a_{1},\cdots,a_{20}$ be the elements of $J_{4}^{\prime}$ defined as follows.
+5. Proof of Theorem 1.1
+In this section, we give our proof of Theorem 1.1.
+First, we will find a fundamental polygon $\widetilde{D}$ on $\mathcal{C}$, equivalently on $\mathbb{H}^{2}$, with respect to the action $\Gamma$ of $PJ_{4}$, and apply Poincaré’s theorem to obtain a presentation of $PJ_{4}$.
+We first prepare the next lemma.
+Lemma 5.1.
+Let $a_{1},\cdots,a_{20}$ be the elements of $J_{4}^{\prime}$ defined as follows.
 $\begin{array}[]{lll}a_{1}:=s_{13}s_{24}s_{12}s_{34}&a_{2}:=s_{13}s_{24}s_{13}s_{24}&a_{3}:=s_{13}s_{34}s_{23}s_{12}\\ a_{4}:=s_{13}s_{34}s_{13}s_{23}&a_{5}:=s_{23}s_{12}s_{23}s_{13}&a_{6}:=s_{23}s_{12}s_{24}s_{12}\\ a_{7}:=s_{23}s_{34}s_{13}s_{34}&a_{8}:=s_{24}s_{34}s_{23}s_{34}&a_{9}:=s_{24}s_{12}s_{24}s_{23}\\ a_{10}:=s_{24}s_{12}s_{23}s_{34}&a_{11}:=s_{24}s_{13}s_{24}s_{13}&a_{12}:=s_{24}s_{23}s_{13}s_{34}\\ a_{13}:=s_{34}s_{23}s_{34}s_{24}&a_{14}:=s_{34}s_{23}s_{12}s_{24}&a_{15}:=s_{34}s_{13}s_{34}s_{23}\\ a_{16}:=s_{34}s_{13}s_{23}s_{24}&a_{17}:=s_{34}s_{12}s_{24}s_{13}&a_{18}:=s_{12}s_{24}s_{12}s_{23}\\ a_{19}:=s_{12}s_{23}s_{34}s_{13}&a_{20}:=s_{12}s_{23}s_{12}s_{13}&\end{array}$
-Then we obtain the next lemma.
-Lemma 4.3.
-The following elements $g_{1},\dots,g_{10}$ and their inverses $g_{1}^{-1},\dots,g_{10}^{-1}$ are contained in $PJ_{4}$.
+Then the following elements $g_{1},\dots,g_{10}$ and their inverses $g_{1}^{-1},\dots,g_{10}^{-1}$ are contained in $PJ_{4}$.
 $\begin{array}[]{ll}g_{1}=a_{1}s_{14}&g_{1}^{-1}=a_{16}s_{14}\\ g_{2}=a_{2}&g_{2}^{-1}=a_{11}\\ g_{3}=a_{3}s_{14}&g_{3}^{-1}=a_{14}s_{14}\\ g_{4}=a_{4}s_{14}&g_{4}^{-1}=a_{9}s_{14}\\ g_{5}=a_{5}&g_{5}^{-1}=a_{20}\\ g_{6}=a_{6}s_{14}&g_{6}^{-1}=a_{15}s_{14}\\ g_{7}=a_{7}s_{14}&g_{7}^{-1}=a_{18}s_{14}\\ g_{8}=a_{8}&g_{8}^{-1}=a_{13}\\ g_{9}=a_{10}s_{14}&g_{9}^{-1}=a_{19}s_{14}\\ g_{10}=a_{12}s_{14}&g_{10}^{-1}=a_{17}s_{14}\end{array}$
-Furthermore, with respect to the action $\Gamma$ on $\mathcal{C}$ in Proposition 4.2, they are all of the elements of $PJ_{4}$ whose translation lengths are at most 4 and actually equal to $4$.
+Furthermore, with respect to the action $\Gamma_{1}$ on $\mathcal{G}$ in Proposition 4.2, they are all of the elements of $PJ_{4}$ whose translation lengths are at most 4 and actually equal to $4$.
 Precisely, if $d(e,g\cdot e)\leq 4$ holds for $g$ in $PJ_{4}$, then $g=g_{i}^{\pm 1}$ and $d(e,g_{i}\cdot e)=4$ for some $1\leq i\leq 10$.
@@
 Thus, by definition, the elements $g_{1},...,g_{10}$ are in $PJ_{4}$.
-Next we show their inverses $g_{1}^{-1},\dots,g_{10}^{-1}$ can be described as in the statement of the lemma.
-For $a_{1},\dots,a_{20}$, the following are obtained by straightforward calculations.
+Next we show their inverses $g_{1}^{-1},\dots,g_{10}^{-1}$ can be described as in the statement of the lemma. For $a_{1},\dots,a_{20}$, the following are obtained by straightforward calculations.
 $\begin{array}[]{llll}a_{1}^{-1}=a_{17}&a_{2}^{-1}=a_{11}&a_{3}^{-1}=a_{19}&a_{4}^{-1}=a_{4}\\ a_{5}^{-1}=a_{20}&a_{6}^{-1}=a_{18}&a_{7}^{-1}=a_{15}&a_{8}^{-1}=a_{13}\\ a_{9}^{-1}=a_{9}&a_{10}^{-1}=a_{14}&a_{12}^{-1}=a_{16}&\end{array}$
-For example,
+For example, we see;
 $\begin{array}[]{ll}a_{1}a_{17}=s_{13}s_{24}s_{12}s_{34}s_{34}s_{12}s_{24}s_{13}=e,\\ a_{17}a_{1}=s_{34}s_{12}s_{24}s_{13}s_{13}s_{24}s_{12}s_{34}=e.\end{array}$
@@
 by simple calculation in the same way as the following.
-$s_{14}a_{1}$ $=s_{14}s_{13}s_{24}s_{12}s_{34}$
-$=s_{24}s_{14}s_{24}s_{12}s_{34}$
-$=s_{24}s_{13}s_{14}s_{12}s_{34}$
-$=s_{24}s_{13}s_{34}s_{14}s_{34}$
-$=s_{24}s_{13}s_{34}s_{12}s_{14}$
-$=s_{24}s_{13}s_{12}s_{34}s_{14}$
-$=s_{24}s_{23}s_{13}s_{34}s_{14}$
-$=a_{12}s_{14}$
+$s_{14}a_{1}$ $=s_{14}s_{13}s_{24}s_{12}s_{34}=s_{24}s_{14}s_{24}s_{12}s_{34}$
+$=s_{24}s_{13}s_{14}s_{12}s_{34}=s_{24}s_{13}s_{34}s_{14}s_{34}$
+$=s_{24}s_{13}s_{34}s_{12}s_{14}=s_{24}s_{13}s_{12}s_{34}s_{14}$
+$=s_{24}s_{23}s_{13}s_{34}s_{14}=a_{12}s_{14}$
 Thus, when we set $g_{1},\dots,g_{20}$ as in the statement, we have
@@
 and so on.
-Finally, with respect to the action $\Gamma$ on $\mathcal{C}$ in Proposition 4.2, we prove that $g_{1},\dots,g_{10}$ and their inverse are all of the elements of $PJ_{4}$ whose translation lengths are equal to $4$, i.e., if $d(e,g\cdot e)\leq 4$ holds for $g$ in $PJ_{4}$, then $g=g_{i}^{\pm 1}$ and $d(e,g_{i}\cdot e)=4$ for some $1\leq i\leq 10$.
-Suppose that an element $g$ of $PJ_{4}$ has translation length at most $4$, i.e. $d(e,g\cdot e)\leq 4$ holds. This means that the element $g\cdot e$ in $J_{4}^{\prime}$ has word length at most 4.
-All the elements of $J_{4}^{\prime}$ whose word lengths are $4$ or less can be listed up by brute force enumerations.
+Finally, with respect to the action $\Gamma_{1}$ on $\mathcal{G}$ in Proposition 4.2, we prove that $g_{1},\dots,g_{10}$ and their inverses are all the elements of $PJ_{4}$ whose translation lengths equal $4$, i.e., if $d(e,g\cdot e)\leq 4$ holds for some $g\in PJ_{4}$, then $g=g_{i}^{\pm 1}$ and $d(e,g_{i}\cdot e)=4$ for some $1\leq i\leq 10$.
+Suppose that an element $g$ of $PJ_{4}$ has translation length at most $4$, i.e., $d(e,g\cdot e)\leq 4$ holds. This means that the element $g\cdot e$ in $J_{4}^{\prime}$ has word length at most $4$.
+All the elements of $J_{4}^{\prime}$ whose word lengths are at most $4$ can be listed by brute-force enumeration.
 See Tables 1 and 2.
-Actually, there are 40 elements of length 3 and 105 elements of length 4.
-If $a:=g\cdot e$ in $J_{4}^{\prime}$, then, by the next claim, it lies on the image of $PJ_{4}$ under the map $\mathfrak{p}$, and so $\pi(a)\in\{e,(14)(23)\}$.
+In fact, there are $40$ elements of length $3$ and $105$ elements of length $4$.
+If $a:=g\cdot e$ in $J_{4}^{\prime}$, then, by the next claim, it lies in the image of $PJ_{4}$ under the map $\mathfrak{p}$, and so $\pi(a)\in\{e,(14)(23)\}$.
 Claim 1.
-Let $\pi$ denotes the natural surjection from $J_{4}$ to $S_{4}$, and $\mathfrak{p}$ the map defined by
-$\mathfrak{p}:PJ_{4}$ $\longrightarrow J_{4}^{\prime}$
-$g$$\longmapsto\begin{cases}g&g\in J_{4}^{\prime}\\ gs_{14}&g\notin J_{4}^{\prime}.\end{cases}$
-Then, the following holds.
-$\mathfrak{p}(PJ_{4})=\left(\pi|_{J_{4}^{\prime}}\right)^{-1}(\{e,(14)(23)\})$
+Let $\pi$ denote the natural surjection from $J_{4}$ to $S_{4}$ and $\mathfrak{p}:PJ_{4}\to J_{4}^{\prime}$ the map defined by
+$g\longmapsto\begin{cases}g&\text{ if }\ g\in J_{4}^{\prime}\\ gs_{14}&\text{ if }\ g\notin J_{4}^{\prime}.\end{cases}$
+Then the following holds:
+$\mathfrak{p}(PJ_{4})=\left(\pi|_{J_{4}^{\prime}}\right)^{-1}(\{e,(14)(23)\}).$
 Proof.
-For any elements $x$ in $\left(\pi|_{J_{4}^{\prime}}\right)^{-1}(\{e,(14)(23)\})$, $\pi(x)=e$ or $(14)(23)$.
-Since $x$ lies on $J_{4}^{\prime}$, $x$ does not contain $s_{14}$ as a word.
-In the former case, $x$ is in $PJ_{4}$ by definition.
-Thus, $\mathfrak{p}(x)=x$.
-In the later case, the following equalities hold.
+For any element $x$ in $\left(\pi|_{J_{4}^{\prime}}\right)^{-1}(\{e,(14)(23)\})$, we have $\pi(x)=e$ or $(14)(23)$.
+Since $x$ lies in $J_{4}^{\prime}$, it does not contain $s_{14}$ as a word.
+In the former case, $x$ is in $PJ_{4}$ by definition, and thus $\mathfrak{p}(x)=x$.
+In the latter case, the following equalities hold.
 $\begin{array}[]{ll}\pi(x)&=(14)(23)\\ \pi(x)\pi(s_{14})&=(14)(23)(14)(23)\\ \pi(xs_{14})&=e\end{array}$
-Therefore, $xs_{14}$ is in $PJ_{4}$ and $xs_{14}$ is not in $J_{4}^{\prime}$.
-It follows $\mathfrak{p}(xs_{14})=xs_{14}^{2}=x$.
-Conversely, each element $x$ in $\mathfrak{p}(PJ_{4})$, $x$ lies on $J_{4}^{\prime}$.
-By the construction of $\mathfrak{p}$, there exists $y$ in $PJ_{4}$ such that $x=y$ or $x=ys_{14}$.
-It follows that $\pi|_{J_{4}^{\prime}}(x)=e$ or $\pi|_{J_{4}^{\prime}}(x)=(14)(23)$.
-Therefore, $x$ is in $\left(\pi|_{J_{4}^{\prime}}\right)^{-1}(\{e,(14)(23)\})$.
-∎
-Among the elements in Tables 1 and 2, by direct checking whether the images of the elements under $\pi$ are $e$ or $(14)(23)$, we have only the elements $a_{1},\cdots,a_{20}$ given in the above.
+Thus, $xs_{14}$ is in $PJ_{4}$ and $xs_{14}$ is not in $J_{4}^{\prime}$. It follows that $\mathfrak{p}(xs_{14})=xs_{14}^{2}=x$.
+Conversely, for each element $x$ in $\mathfrak{p}(PJ_{4})$, we have $x\in J_{4}^{\prime}$.
+By the construction of $\mathfrak{p}$, there exists $y\in PJ_{4}$ such that $x=y$ or $x=ys_{14}$.
+It follows that
+$\pi|_{J_{4}^{\prime}}(x)=e\quad\text{or}\quad\pi|_{J_{4}^{\prime}}(x)=(14)(23).$
+Therefore, $x\in\left(\pi|_{J_{4}^{\prime}}\right)^{-1}(\{e,(14)(23)\})$. ∎
+Among the elements in Tables 1 and 2, by directly checking whether the images of the elements under $\pi$ are $e$ or $(14)(23)$, we obtain only the elements $a_{1},\cdots,a_{20}$ as given above.
 (Remark that $a_{16}=s_{34}s_{13}s_{23}s_{24}$ appears as $s_{34}s_{13}s_{24}s_{34}$ in Table 2.)
-Consequently, all of elements of $PJ_{4}$ whose translation lengths are at most $4$ are shown to be $g_{1},...,g_{10}$ and their inverse elements $g_{1}^{-1},\dots,g_{10}^{-1}$. ∎
+Consequently, all elements of $PJ_{4}$ whose translation lengths are at most $4$ are shown to be $g_{1},\dots,g_{10}$ and their inverse elements $g_{1}^{-1},\dots,g_{10}^{-1}$. ∎
 $\begin{array}[]{cccccc}s_{13}s_{23}s_{34}&s_{13}s_{23}s_{24}&s_{13}s_{24}s_{12}&s_{13}s_{24}s_{13}&s_{13}s_{24}s_{23}&s_{13}s_{34}s_{23}\\ s_{13}s_{34}s_{13}&s_{13}s_{34}s_{12}&s_{23}s_{13}s_{24}&s_{23}s_{13}s_{23}&s_{23}s_{12}s_{23}&s_{23}s_{12}s_{24}\\ s_{23}s_{12}s_{34}&s_{23}s_{34}s_{13}&s_{23}s_{34}s_{23}&s_{23}s_{34}s_{24}&s_{24}s_{34}s_{13}&s_{24}s_{34}s_{12}\\ s_{24}s_{12}s_{24}&s_{24}s_{12}s_{23}&s_{24}s_{12}s_{13}&s_{24}s_{13}s_{24}&s_{24}s_{13}s_{34}&s_{24}s_{13}s_{12}\\ s_{34}s_{24}s_{12}&s_{34}s_{24}s_{34}&s_{34}s_{23}s_{34}&s_{34}s_{23}s_{12}&s_{34}s_{23}s_{13}&s_{34}s_{13}s_{34}\\ s_{34}s_{13}s_{24}&s_{34}s_{13}s_{23}&s_{12}s_{34}s_{23}&s_{12}s_{34}s_{24}&s_{12}s_{24}s_{13}&s_{12}s_{24}s_{12}\\ s_{12}s_{24}s_{34}&s_{12}s_{23}s_{34}&s_{12}s_{23}s_{12}&s_{12}s_{23}s_{13}&&\end{array}$
@@
 Table 2. Elements of $J_{4}^{\prime}$ of length 4
-Let $\widetilde{D}$ be the polygon on $\mathcal{C}$ defined by following, which is is visualized in Figure 4.
-$\displaystyle\widetilde{D}:=\bigcap_{i=e}^{20}\left\{x\in\mathcal{C}\mid\ d(e,x)\leq d(x,a_{i})\right\}$
+Let $\widetilde{D}$ be the polygon on $\mathcal{C}$ defined as follows, which is visualized in Figure 4.
+$\displaystyle\widetilde{D}:=\bigcap_{i=1}^{20}\left\{x\in\mathcal{C}\mid\ d(e,x)\leq d(x,a_{i})\right\}$
 […]
 Figure 4.
-From Figure 4, we see the following.
+From Figure 4, the following can be verified.
 Lemma 5.2.
-The word length of the vertices $\mathrm{int}\widetilde{D}$ are at most $3$.
+With respect to the action $\Gamma_{1}$ on $\mathcal{G}$, the word lengths of the vertices in $\mathrm{int}\,\widetilde{D}$ are at most $3$.
 The vertices on $\partial\widetilde{D}$ whose word length is $3$ are only $s_{13}s_{24}s_{23}$, $s_{23}s_{34}s_{12}$, $s_{24}s_{13}s_{23}$, $s_{34}s_{13}s_{12}$, $s_{12}s_{24}s_{34}$.
-We denote the five vertices above by $v_{1}$, …, $v_{5}$, respectively.
+∎
+We denote the five vertices above by $v_{1}$, $\dots$, $v_{5}$, respectively.
 We remark that there are ten sides of $\widetilde{D}$ which are not edges in the Cayley complex $\mathcal{C}$.
 These sides are the diagonals of some 2-cells (quadrangles) in $\mathcal{C}$.
-The interior angles around $v_{1}$, …, $v_{5}$ are $\frac{2\pi}{5}$.
-The interior angles around $s_{13}s_{23}$, $s_{13}s_{12}$, $s_{24}s_{34}$, $s_{24}s_{23}$, $s_{34}s_{12}$ are $\frac{4\pi}{5}$.
-The interior angles around $s_{12}s_{23}$, $s_{13}s_{24}$, $s_{13}s_{34}$, $s_{23}s_{12}$, $s_{23}s_{34}$, $s_{24}s_{12}$, $s_{24}s_{13}$, $s_{34}s_{23}$, $s_{34}s_{13}$, $s_{12}s_{24}$ are $\frac{3\pi}{5}$.
+The interior angles at $v_{1}$, $\dots$, $v_{5}$ are $\frac{2\pi}{5}$.
+The interior angles at the vertices $s_{13}s_{23}$, $s_{13}s_{12}$, $s_{24}s_{34}$, $s_{24}s_{23}$, $s_{34}s_{12}$ on $\partial\widetilde{D}$ are $\frac{4\pi}{5}$.
+The interior angles at $s_{12}s_{23}$, $s_{13}s_{24}$, $s_{13}s_{34}$, $s_{23}s_{12}$, $s_{23}s_{34}$, $s_{24}s_{12}$, $s_{24}s_{13}$, $s_{34}s_{23}$, $s_{34}s_{13}$, $s_{12}s_{24}$ on $\partial\widetilde{D}$ are $\frac{3\pi}{5}$.
 Lemma 5.3.
-The polygon $\widetilde{D}$ is the Dirichlet polygon with center $e$ for $PJ_{4}$ as a subgroup of $\mathrm{Isom}\;\mathcal{C}$.
+The subset $\widetilde{D}_{1}:=\mathrm{int}(\widetilde{D}\cap\mathcal{G})$ of $\mathcal{G}$ is the Dirichlet domain with center $e$ in $\mathcal{G}$ with respect to the action $\Gamma_{1}$ of $PJ_{4}$.
 Proof.
-Since $PJ_{4}$ is nontrivial, we consider
+Since $PJ_{4}$ is nontrivial, by definition, we consider
 $D(e)=\bigcap_{g\neq e}H_{g}(e).$
-Here $H_{g}(e)$ denotes the half space defined by $\{x\in\mathcal{C}\mid d(x,e)<d(x,g\cdot e)\}$ for $g$ in $PJ_{4}$.
-[…]
-Let $g$ be a non-trivial element of $PJ_{4}$.
-By Lemma 4.3, if $g$ satisfies that $d(e,g\cdot e)$ is $4$ or less, then it is equal to $4$ and $g=g_{i}^{\pm 1}$ for some $i$.
-Therefore, $H_{g}(e)$ is equal to $H_{g_{i}^{\pm 1}}(e)$ for some $i$ when $d(e,g\cdot e)$ is 4 or less.
-If $d(e,g\cdot e)$ is $6$ or more, we obtain the following.
-$\{x\mid d(e,x)<3\}\subset H_{g}(e)$
-[…]
-It remains to show that $d(e,g\cdot e)\neq 5$ on $\mathcal{C}$ for any $g$ in $PJ_{4}$.
-Recall that the image of the standard generators $s_{12}$, $s_{23}$, $s_{34}$, $s_{13}$, $s_{24}$ of $J_{4}^{\prime}$ under $\pi$ are the transpositions $(12)$, $(23)$, $(34)$, $(13)$, $(24)$ in $S_{4}$, and also that $\pi(s_{14})=(14)(23)$.
+Here, $H_{g}(e)$ denotes the half-space defined by $\{x\in\mathcal{G}\mid d(x,e)<d(x,g\cdot e)\}$ for $g$ in $PJ_{4}$.
+We prove that $\widetilde{D}_{1}$ coincides with $D(e)$.
+By the definition of the Dirichlet domain, $\widetilde{D}_{1}$ contains $D(e)$.
+Thus, it suffices to show the converse, i.e., that $H_{g}(e)$ contains $\widetilde{D}_{1}$ for any nontrivial element $g$ in $PJ_{4}$.
+Let $g$ be a nontrivial element of $PJ_{4}$.
+By Lemma 5.1, if $g$ satisfies $d(e,g\cdot e)\leq 4$, then $d(e,g\cdot e)=4$ and $g=g_{i}^{\pm 1}$ for some $i$.
+Therefore, $H_{g}(e)$ is equal to $H_{g_{i}^{\pm 1}}(e)$ for some $i$ when $d(e,g\cdot e)\leq 4$.
+If $d(e,g\cdot e)\geq 6$, we obtain the following:
+$\{x\mid d(e,x)<3\}\subset H_{g}(e).$
+Since the word lengths of the vertices in $\widetilde{D}_{1}$ are less than $3$, we have $d(e,x)<3$ for any $x\in\widetilde{D}_{1}$. Thus,
+$\bigcap\{H_{g}(e)\mid d(e,g\cdot e)\geq 6\}\supset\widetilde{D}_{1}.$
+It remains to show that $d(e,g\cdot e)\neq 5$ on $\mathcal{G}$ for any $g\in PJ_{4}$.
+Recall that the images of the standard generators $s_{12}$, $s_{23}$, $s_{34}$, $s_{13}$, $s_{24}$ of $J_{4}^{\prime}$ under $\pi$ are the transpositions $(12)$, $(23)$, $(34)$, $(13)$, $(24)$ in $S_{4}$, and also that $\pi(s_{14})=(14)(23)$.
 See Section 2.
-Let $g$ be an element in $PJ_{4}$. It means that $\pi(g)=e$ in $S_{4}$.
+Let $g$ be an element in $PJ_{4}$. This means that $\pi(g)=e$ in $S_{4}$.
 If $g$ is contained in $J_{4}^{\prime}$, then $g\cdot e=g$ is in $J_{4}^{\prime}$.
-Thus $d(e,g\cdot e)=d(e,g)=5$ implies $g$ has word length 5 with respect to the generators $s_{12}$, $s_{23}$, $s_{34}$, $s_{13}$, $s_{24}$ in $J_{4}^{\prime}$.
-It would imply that $\pi(g)$ is an odd permutation, contradicting that $\pi(g)=e$.
+Thus $d(e,g\cdot e)=d(e,g)=5$ implies that $g$ has word length 5 with respect to the generators $s_{12}$, $s_{23}$, $s_{34}$, $s_{13}$, $s_{24}$ in $J_{4}^{\prime}$.
+This would imply that $\pi(g)$ is an odd permutation, contradicting $\pi(g)=e$.
 If $g$ is not contained in $J_{4}^{\prime}$, then $g\cdot e=gs_{14}$ is in $J_{4}^{\prime}$ by definition of the action $\Gamma$.
-Thus $d(e,g\cdot e)=d(e,gs_{14})=5$ implies $gs_{14}$ has word length 5 with respect to the generators $s_{12}$, $s_{23}$, $s_{34}$, $s_{13}$, $s_{24}$ in $J_{4}^{\prime}$, that is, $\pi(gs_{14})$ must be an odd permutation.
-Since $\pi(s_{14})=(14)(23)$, it would imply that $\pi(g)$ is an odd permutation, contradicting that $\pi(g)=e$.
-As a result, $D(e)$ contains the interior of $\widetilde{D}$, and it concludes that the polygon $\widetilde{D}$ is the Dirichlet polygon with center $e$ for $PJ_{4}$ as a subgroup of $\mathrm{Isom}\;\mathcal{C}$.
+Thus $d(e,g\cdot e)=d(e,gs_{14})=5$ implies $gs_{14}$ has word length 5 with respect to the generators $s_{12}$, $s_{23}$, $s_{34}$, $s_{13}$, $s_{24}$ in $J_{4}^{\prime}$, and $\pi(gs_{14})$ must be an odd permutation.
+Since $\pi(s_{14})=(14)(23)$, it implies that $\pi(g)$ is an odd permutation, contradicting $\pi(g)=e$.
+As a result, $D(e)$ contains the interior of $\widetilde{D}_{1}$.
+It concludes that $\widetilde{D}_{1}=\mathrm{int}(\widetilde{D}\cap\mathcal{G})$ is the Dirichlet domain with center $e$ on $\mathcal{G}$ with respect to the action $\Gamma_{1}$ of $PJ_{4}$.
 ∎
@@
 Proof of Theorem 1.1.
-By Lemma 4.5, the interior of $\widetilde{D}$ is the Dirichlet domain for the action $\Gamma$ by $PJ_{4}$.
-Since the group which acts on a topological space freely and cocompactly is discontinuous, togethter with Proposition 4.2, $PJ_{4}$ is shown to be discontinuous.
-By Proposition 4.1, $\mathcal{C}$ is geodesically conected, geodesically complete, and finitely compact.
+By Lemma 5.3, it follows that the interior of $\widetilde{D}$ is the Dirichlet domain on $\mathcal{C}$, and thus $\widetilde{D}$ is a Dirichlet polygon on $\mathcal{C}$ with respect to the action $\Gamma$ of $PJ_{4}$ on $\mathcal{C}$.
+This is because the action $\Gamma$ on $\mathcal{C}$ is a natural extension of $\Gamma_{1}$ on $\mathcal{G}=\mathcal{C}^{(1)}$, which is an isometric action on $\mathcal{C}$, and $\widetilde{D}$ is the convex hull in $\mathcal{C}$ of $\widetilde{D}\cap\mathcal{G}$.
+Since a group acting freely and cocompactly on a topological space is discontinuous, together with Proposition 4.2, it follows that $PJ_{4}$ is discontinuous.
+By Proposition 4.1, $\mathcal{C}$ is geodesically connected, geodesically complete, and finitely compact.
 Thus, applying Proposition 3.1 to the interior of $\widetilde{D}$, we obtain that $\widetilde{D}$ is a locally finite fundamental polygon for $PJ_{4}$.
 By Proposition 3.2, $PJ_{4}$ is generated by the elements $g_{1},\dots,g_{10}$ and their inverses.
-Next, we will obtain the relations with respect to $g_{1},...,g_{10}$ by applying Proposition 3.3.
-The generators $g_{1},...,g_{10}$ of $PJ_{4}$ give identifications on (the sides of) the polygon $\widetilde{D}$.
+Next, we will obtain the relations for $g_{1},\dots,g_{10}$ by applying Proposition 3.3.
+The generators $g_{1},\dots,g_{10}$ of $PJ_{4}$ define identifications on the sides of the polygon $\widetilde{D}$.
 For example, by
@@
 $\langle g_{2},g_{4},g_{8},g_{9},g_{10}\mid g_{2}g_{9}g_{10}^{-1}g_{8}^{-1}g_{4}g_{9}g_{2}g_{10}g_{8}^{-1}g_{4}^{-1}=e\rangle$
-We remark that, by calculating the Euler characteristic, we see that the surface obtained as $\widetilde{D}^{*}$ is homeomorphic to the connected sum of the five projective plane.
+We remark that, by calculating the Euler characteristic, we see that the surface obtained from $\widetilde{D}$ by gluing its sides via $g_{j}$’s is homeomorphic to the connected sum of the five projective plane.
 6. Proof of corollaries
@@
 By virtue of geometric descriptions, we could find the correspondence between $\alpha_{i}$ and $g_{j}$ in the proof above as follows.
-The painted regions in Figure 5 give five Möbius bands embedded in $\mathcal{C}/PJ_{4}$.
-After removing them, we can rebuild the remaining parts to obtain a sphere with five holes.
+The painted regions in Figure 5 represent five Möbius bands embedded in $\mathcal{C}/PJ_{4}$.
+After removing them, the remaining parts can be reassembled to form a sphere with five holes.
 See Figure 6.
-Then we can confirm that $\mathcal{C}/PJ_{4}$ is homeomorphic to the connected sum #$\#_{5}\mathbb{RP}^{2}$ of five real projective planes.
-Setting the closed curve $\alpha_{2}$ in #$\#_{5}\mathbb{RP}^{2}$ shown in Figure 6, we obtain a correspondence $\alpha_{2}\longmapsto g_{8}^{-1}g_{4}g_{9}g_{2}g_{10}\in PJ_{4}$ by considering a lift of $\alpha_{2}$ in $\mathcal{C}$.
-In the same way, we can obtain the corresponding elements in $PJ_{4}$ for $\alpha_{1}$, $\alpha_{3}$, $\alpha_{4}$, $\alpha_{5}$.
+Then, we can confirm that $\mathcal{C}/PJ_{4}$ is homeomorphic to the connected sum #$\#_{5}\mathbb{RP}^{2}$ of five real projective planes.
+By setting the closed curve $\alpha_{2}$ in #$\#_{5}\mathbb{RP}^{2}$ shown in Figure 6, we obtain a correspondence
+$\alpha_{2}\longmapsto g_{8}^{-1}g_{4}g_{9}g_{2}g_{10}\in PJ_{4}$
+by considering a lift of $\alpha_{2}$ in $\mathcal{C}$. Similarly, we obtain the corresponding elements in $PJ_{4}$ for $\alpha_{1}$, $\alpha_{3}$, $\alpha_{4}$, and $\alpha_{5}$.
 […]